### v2

# Convergence analysis of Lawson’s iteration for the polynomial and rational minimax approximations

Abstract
Lawson’s iteration is a classical and effective method for solving the linear (polynomial) minimax approximation in the complex plane. Extension of Lawson’s iteration for the rational minimax approximation with both computationally high efficiency and theoretical guarantee is challenging. A recent work [L.-H. Zhang, L. Yang, W. H. Yang and Y.-N. Zhang, A convex dual programming for the rational minimax approximation and Lawson’s iteration, 2023, https://arxiv.org/pdf/2308.06991v1] reveals that Lawson’s iteration can be viewed as a method for solving the dual problem of the original rational minimax approximation, and a new type of Lawson’s iteration was proposed. Such a dual problem is guaranteed to obtain the original minimax solution under Ruttan’s sufficient condition, and numerically, the proposed Lawson’s iteration was observed to converge monotonically with respect to the dual objective function. In this paper, we perform theoretical convergence analysis for Lawson’s iteration for both the linear and rational minimax approximations. In particular, we show that
  • (i)
    for the linear minimax approximation, the near-optimal Lawson exponent $\beta$ in Lawson’s iteration is $\beta=1$, and
  • (ii)
    for the rational minimax approximation, the proposed Lawson’s iteration converges monotonically with respect to the dual objective function for any sufficiently small $\beta>0$, and the limit approximant fulfills the complementary slackness: any node associated with positive weight either is an interpolation point or has a constant error.
Key words. Rational minimax approximation, Lawson algorithm, Ruttan’s optimality condition, Dual programming, Convergence analysis
AMS subject classifications. 41A50, 41A20, 65D15, 90C46
1 Introduction
Computing the polynomial and/or rational minimax (also known as Chebyshev or best) approximations of a function $f$ over a given set $\Omega$ in the complex plane $\mathbb{C}$ is a classical problem in approximation theory [34]. In practical applications, rational approximations are useful in many areas, including conformal mapping [14, 35], model order reduction [15, 2, 7], and signal processing [8, 17, 36, 37]. See [25] for some recent applications. In some cases, $\Omega$ is a compact set in $\mathbb{C}$ and $f$ is a continuous complex-valued function on $\Omega$. Also, in many real-world applications, for example, in signal processing and microwave duplexer filtering [1, 12, 13, 33], only discretized sample data are available; even for a continuum domain $\Omega$ enclosed by a simple Jordan curve in which $f$ is analytic, by the maximum modulus principle, we can first sample $f$ on the boundary of $\Omega$, and then compute the rational/polynomial minimax approximant of $f$ through solving a discrete rational/polynomial minimax problem. For these cases, denote by $\{(x_{j},f_{j})\}_{j=1}^{m}$ the sampled data from $f_{j}=f(x_{j})\in\mathbb{C}$ ($x_{j}\in\Omega$) over distinct nodes ${\cal X}=\{x_{j}\}_{j=1}^{m}$, and denote by $\mathbb{P}_{n}$ the set of complex polynomials with degree less than or equal to $n$, we consider the following discrete rational approximation
$\inf_{\xi=p/q\in\mathscr{R}_{(n_{1},n_{2})}}\|\boldsymbol{f}-\xi(\boldsymbol{x})\|_{\infty},$ (1.1)
where $\mathscr{R}_{(n_{1},n_{2})}:=\{{p}/{q}|p\in\mathbb{P}_{n_{1}},~{}0\not\equiv q\in\mathbb{P}_{n_{2}}\}$, $\boldsymbol{f}=[f_{1},\dots,f_{m}]^{\operatorname{T}}\in\mathbb{C}^{m}~{}(n_{1}+n_{2}+2\leq m)$, $\boldsymbol{x}=[x_{1},\dots,x_{m}]^{\operatorname{T}}\in\mathbb{C}^{m}$, $\xi(\boldsymbol{x})=[\xi(x_{1}),\dots,\xi(x_{m})]^{\operatorname{T}}\in\mathbb{C}^{m}$, and
$\|\boldsymbol{f}-\xi(\boldsymbol{x})\|_{\infty}:=\max_{1\leq j\leq m}\left|f_{j}-\frac{p(x_{j})}{q(x_{j})}\right|.$
In case when the infimum of (1.1) is attainable, we call the function $\xi^{*}={p^{*}}/{q^{*}}\in\mathscr{R}_{(n_{1},n_{2})}$ from
${p^{*}}/{q^{*}}\in\arg\min_{\xi=p/q\in\mathscr{R}_{(n_{1},n_{2})}}\|\boldsymbol{f}-\xi(\boldsymbol{x})\|_{\infty},$ (1.2)
the rational minimax approximant [34] of $f(x)$ over ${\cal X}$.
In general, computing the discrete rational minimax approximation is much more challenging than the polynomial (i.e., $n_{2}=0$) minimax problem. Indeed, for the polynomial case, it is guaranteed that there is a unique minimax approximant which can be characterized by a necessary and sufficient condition (e.g., [28] and [38, Theorem 2.1]), while for the rational case, the infimum of (1.1) may not be achievable, and even it does, there may be multiple minimax approximants [24, 34]. Furthermore, local but non-global minimax approximants may exist [24, 34]. Necessary conditions for the rational minimax approximant have been developed in, e.g., [10, 16, 29, 30, 32, 38, 39, 40], and Ruttan [29, Theorems 2.1] contributes a sufficient condition.
Lawson’s iteration [21] is a classical and effective method for computing the discrete linear (polynomial) minimax approximant. The idea of Lawson is to approximate the minimax polynomial $p^{*}$ by a sequence of polynomials $\{p^{(k)}\}$, each as the solution of the weighted least-squares problem:
$p^{(k)}=\operatorname*{argmin}_{p\in\mathbb{P}_{n}}\sum_{j=1}^{m}w_{j}^{(k)}|f_{j}-p(x_{j})|^{2},$
where $\boldsymbol{w}^{(k)}=[w_{1}^{(k)},\dots,w_{m}^{(k)}]^{\operatorname{T}}\in{\cal S}$ is the weight vector at the $k$th iteration in the probability simplex:
${\cal S}:=\{\boldsymbol{w}=[w_{1},\dots,w_{m}]^{\operatorname{T}}\in\mathbb{R}^{m}:\boldsymbol{w}\geq 0~{}{\rm and}~{}\boldsymbol{w}^{\operatorname{T}}\boldsymbol{e}=1\},~{}~{}\boldsymbol{e}=[1,1,\dots,1]^{\operatorname{T}}.$
A Lawson’s step at the $k$the iteration updates the weight vector as
$w_{j}^{(k+1)}=\frac{w_{j}^{(k)}\left|f_{j}-p^{(k)}(x_{j})\right|^{\beta}}{\sum_{i}w_{i}^{(k)}\left|f_{i}-p^{(k)}(x_{i})\right|^{\beta}},~{}~{}\forall j,$ (1.3)
where $\beta>0$ is the so-called Lawson exponent and is originally set as $\beta=1$ ([21, 5]). Due to its relation with weighted least-squares problems, Lawson’s iteration is an iteratively reweighted least-squares (IRLS) iteration. Convergence analysis and some variants have been discussed (e.g., [5, 3, 6, 18, 22, 31, 41]).
However, extension of Lawson’s iteration for the rational minimax approximation problem with both computationally high efficiency and theoretical guarantee is nontrivial and challenging. Particularly, for computing the minimax rational $\xi^{*}=p^{*}/q^{*}$ of (1.2), two difficulties related with a basic Lawson’s step (1.3) are: how to define a suitable approximation $\xi^{(k)}=p^{(k)}/q^{(k)}$ associated with the current weight vector $\boldsymbol{w}^{(k)}$, and how to choose a suitable Lawson exponent $\beta$ for convergence?
Some versions of Lawson’s iteration have been discussed for the rational minimax approximation (1.1). For example, the Loeb algorithm [22] (the same method was also proposed in [31] known as the SK iteration by Sanathanan and Koerner) uses the reciprocal of the current denominator $q^{(k)}(x_{j})$ as weights and compute the approximation $\xi^{(k)}=p^{(k)}/q^{(k)}$ from a weighted linearization associated with the current weights. A recent work [18] further improves the basic SK iteration by proposing a stabilized SK iteration. Another remarkable work on the rational approximation is the adaptive Antoulas-Anderson (AAA) [23] and its extension, AAA-Lawson algorithm [11, 24, 25]. AAA represents the rational approximation in barycentric form and selects the associated support points iteratively in an adaptive way for stability; in AAA-Lawson [11, 24, 25], the algorithm further introduces weights updated according to [11, Equ. (8.5)]
$w_{j}^{(k+1)}=\frac{w_{j}^{(k)}\left|f_{j}-\xi_{j}^{(k)}(x_{j})\right|^{\beta}}{\sum_{i}w_{i}^{(k)}\left|f_{i}-\xi_{i}^{(k)}(x_{j})\right|^{\beta}},~{}~{}\forall j,$ (1.4)
where $\xi_{j}^{(k)}$ is from a weighted linearization [11, Equ. (8.4)] associated with the current weight vector $\boldsymbol{w}^{(k)}$. Other versions of Lawson’s iteration can be found in [3, 6]. However, to our best knowledge, no convergence guarantee has been established for these versions of Lawson’s iteration in theory, and in some cases, the computed rational approximants can be local best or near-best [11, 24]. Indeed, as remarked for AAA-Lawson in [11] that “its convergence is far from understood, and even when it does converge, the rate is slow (linear at best)”, and “convergence analysis appears to be highly nontrivial”.
In this paper, we shall establish the convergence of a version of Lawson’s iteration (Algorithm 1) proposed recently [43]. In [43], it reveals that Lawson’s iteration can be viewed as a method for solving the dual problem $\max_{\boldsymbol{w}\in{\cal S}}d(\boldsymbol{w})$ of the original rational minimax approximation. The weight $w_{j}$ is the corresponding dual variable associated with the node $x_{j}$. Such a dual problem is guaranteed to obtain the original minimax solution $\xi^{*}$ under Ruttan’s sufficient condition ([29, Theorems 2.1]; see also [19, Theorem 2] and [32, Theorem 3]); moreover, numerically, the proposed Lawson’s iteration was observed to converge monotonically [43] with respect to the dual objective function $d(\boldsymbol{w})$. The framework for handling the rational minimax approximation (1.1) in [43] can be well described by Figure 1.1 (see [43, Figure 1.1]). It should be pointed out that this Lawson’s iteration reduces to the classical Lawson’s iteration [21] for the linear minimax approximation if $n_{2}=0$.
[…]
Figure 1.1: Framework [43, Figure 1.1] of the dual programming and Lawson’s iteration for the rational minimax approximation of (1.2)
For this version of Lawson’s iteration (Algorithm 1), we shall contribute theoretical convergence analysis for both the linear and rational minimax approximations. In particular, we show that
  • (i)
    for the linear minimax approximation, $\beta=1$ is a near-optimal Lawson exponent in Lawson’s iteration (1.3), and
  • (ii)
    for the rational minimax approximation, the proposed Lawson’s iteration converges monotonically with respect to the dual objective function $d(\boldsymbol{w})$ for any sufficiently small $\beta>0$ in (1.4), and moreover, the convergent pair $(\boldsymbol{w},\xi)$ fulfills the following complementary slackness:
    $w_{j}r_{j}\left(|f_{j}-\xi(x_{j})|-c\right)=0,\quad\forall j=1,2,\dots,m,$
    for some constant $c$. The complementary slackness means that any node associated with positive weight either is an interpolation point or has a constant error.
These theoretical findings lay a solid ground for this Lawson’s iteration (Algorithm 1).
We organize the paper as follows. In section 2, we first introduce the dual programming associated with a linearization of the rational minimax problem (1.1) in the Lagrange duality [43]. These results are mainly from [43]. We shall discuss the strong duality and its relation with Ruttan’s sufficient condition; also optimality condition for computing the dual objective function and Lawson’s iteration [43] (Algorithm 1) for the dual programming will be presented. In section 3, we shall establish an important lower bound for the dual objective function value $d(\boldsymbol{w}^{(k+1)})$ based on the information at the $k$th iteration. Relying on this lower bound, we prove in section 4 that for the linear minimax approximation, there is a $\beta_{0}>1$ so that for any $\beta\in(0,\beta_{0})$, Lawson’s iteration converges monotonically, and $\beta=1$ is the near-optimal Lawson exponent. In section 5, we will use the lower bound to show that monotonic convergence of Lawson’s iteration occurs generically for any sufficiently small $\beta>0$, and furthermore, we shall prove the complementary slackness whenever $d(\boldsymbol{w}^{(k+1)})=d(\boldsymbol{w}^{(k)})$. Finally concluding remarks are drawn in section 6.
Notation. We follow the notation in [43] in this paper. The imaginary unit is ${\tt i}=\sqrt{-1}$, and for $\mu=\mu^{\tt r}+{\tt i}\mu^{\tt i}\in\mathbb{C}$, we denote its modulus $|\mu|=\sqrt{(\mu^{\tt r})^{2}+(\mu^{\tt i})^{2}}$ and its conjugate $\bar{\mu}=\mu^{\tt r}-{\tt i}\mu^{\tt i}$, where ${\rm Re}(\mu)=\mu^{\tt r}\in\mathbb{R}$ and ${\rm Im}(\mu)=\mu^{\tt i}\in\mathbb{R}$ are the real and imaginary part of $\mu$, respectively. Bold lower case letters are used to represent column vectors, and ${\mathbb{C}}^{n\times m}$ (resp. ${\mathbb{R}}^{n\times m}$) stands for the set of all $n\times m$ complex (resp. real) matrices, with the identity $I_{n}\equiv[\boldsymbol{e}_{1},\boldsymbol{e}_{2},\dots,\boldsymbol{e}_{n}]\in\mathbb{R}^{n\times n}$, where $\boldsymbol{e}_{i}$ is its $i$th column with $i\in[n]:=\{1,2,\dots,n\}$. For a vector $\boldsymbol{x}\in\mathbb{C}^{n}$, ${\rm diag}(\boldsymbol{x})={\rm diag}(x_{1},\dots,x_{n})$ is the diagonal matrix, and $\|\boldsymbol{x}\|_{\alpha}=(\sum_{j=1}^{n}|x_{j}|^{\alpha})^{\frac{1}{\alpha}}$ is the vector $\alpha$-norm ($\alpha\geq 1$) of $\boldsymbol{x}$. For $\boldsymbol{x},\boldsymbol{y}\in\mathbb{C}^{n}$ with $y_{j}\neq 0,~{}1\leq j\leq n$, we define
$\boldsymbol{x}./\boldsymbol{y}=[x_{1}/y_{1},\dots,x_{n}/y_{n}]^{\operatorname{T}}.$
For a matrix $A\in\mathbb{C}^{m\times n}$, ${\rm span}(A)$ represents the column space of $A$; $A^{\operatorname{H}}$ (resp. $A^{\operatorname{T}}$) and $A^{{\dagger}}$ are the conjugate transpose (resp. transpose) and the Moore-Penrose inverse of $A$, respectively. We also adopt MATLAB-like convention to represent the sub-matrix $A({\cal I}_{1},{\cal I}_{2})$ of $A$, consisting of intersections of rows and columns indexed by ${\cal I}_{1}\subseteq[m]$ and ${\cal I}_{2}\subseteq[n]$, respectively.
2 A dual programming and Lawson’s iteration
First, let $\mathbb{P}_{n_{1}}={\rm span}(\psi_{0}(x),\dots,\psi_{n_{1}}(x))$ and $\mathbb{P}_{n_{2}}={\rm span}(\phi_{0}(x),\dots,\phi_{n_{2}}(x))$ be the chosen bases for the numerator and denominator polynomial spaces, respectively, and we write $p/q\in\mathscr{R}_{(n_{1},n_{2})}$ as
[…]
For the given ${\cal X}=\{x_{j}\}_{j=1}^{m}$ with $|{\cal X}|=m$, we have the basis matrix for $p\in\mathbb{P}_{n_{1}}$:
$\Psi=\Psi(x_{1},\dots,x_{m};n_{1}):=\left[\begin{array}[]{cccc}\psi_{0}(x_{1})&\psi_{1}(x_{1})&\cdots&\psi_{n_{1}}(x_{1})\\ \psi_{0}(x_{2})&\psi_{1}(x_{2})&\cdots&\psi_{n_{1}}(x_{2})\\ \vdots&\cdots&\cdots&\vdots\\ \psi_{0}(x_{m})&\psi_{1}(x_{m})&\cdots&\psi_{n_{1}}(x_{m})\end{array}\right],~{}\Psi_{i,j}=\psi_{j-1}(x_{i}),$
and analogously, we have $\Phi=\Phi(x_{1},\dots,x_{m};n_{2})=[\phi_{j-1}(x_{i})]\in\mathbb{C}^{m\times(n_{2}+1)}$.
For a given $\xi(x)=p(x)/q(x)\in\mathscr{R}_{(n_{1},n_{2})}$, if $|\xi(x)|$ is bounded for any $x\in{\cal X}$, then it is easy to see that $q(x)\neq 0$. We define the maximum error
$e(\xi):=\max_{x\in{\cal X}}|f(x)-\xi(x)|=\|\boldsymbol{f}-\xi(\boldsymbol{x})\|_{\infty}.$ (2.1)
The defect of an irreducible $\xi(x)=p(x)/q(x)\in\mathscr{R}_{(n_{1},n_{2})}$ is
$\upsilon(p,q):=\min(n_{1}-\deg(p),n_{2}-\deg(q)),$ (2.2)
where $\deg(p)$ and $\deg(q)$ are the degrees of $p$ and $q$, respectively. When $\upsilon(p,q)=0$, we say $\xi(x)=p(x)/q(x)$ is non-degenerate. As a necessary condition for the irreducible minimax rational approximant $\xi^{*}=p^{*}/q^{*}$ of (1.2), the following result states that the number of reference points (aks the extreme points), i.e., nodes $x_{j}\in{\cal X}$ that achieves
$\left|f_{j}-\frac{p^{*}(x_{j})}{q^{*}(x_{j})}\right|=e(\xi^{*}),$
is at least $n_{1}+n_{2}+2-\upsilon(p^{*},q^{*})$.
Theorem 2.1.
([16, Theorem 2.5]) Given $m\geq{n_{1}}+n_{2}+2$ distinct nodes ${\cal X}=\{x_{j}\}_{j=1}^{m}$ on $\Omega$, suppose $\xi^{*}=p^{*}/q^{*}\in\mathscr{R}_{(n_{1},n_{2})}$ is an irreducible rational polynomial and denote the extremal set ${\cal X}_{e}(\xi^{*})\subseteq{\cal X}$ by
${\cal X}_{e}(\xi^{*}):=\left\{x_{j}\in{\cal X}:\left|f_{j}-\frac{p^{*}(x_{j})}{q^{*}(x_{j})}\right|=e(\xi^{*})\right\}.$ (2.3)
If $\xi^{*}$ is a solution to (1.2) with $\eta_{\infty}=\|\boldsymbol{f}-\xi^{*}(\boldsymbol{x})\|_{\infty},$ then the cardinality $|{\cal X}_{e}(\xi^{*})|\geq n_{1}+n_{2}+2-\upsilon(p^{*},q^{*})$; that is, ${\cal X}_{e}(\xi^{*})$ contains at least $n_{1}+n_{2}+2-\upsilon(p^{*},q^{*})$ nodes.
In the following discussion, we assume, without loss of generality, that $\eta_{\infty}>0$, i.e., $\mathscr{R}_{(n_{1},n_{2})}$ does not contain an interpolation for $\boldsymbol{f}$.
2.1 A linearization
In [43], by introducing a real variable $\eta$, the original minimax problem (1.1) is transformed into the following linearization
$\inf_{\eta\in\mathbb{R},~{}p\in\mathbb{P}_{n_{1}},~{}q\in\mathbb{P}_{n_{2}}\setminus\{0\}}\eta$
$s.t.,~{}$ $|f_{j}q(x_{j})-p(x_{j})|^{\alpha}\leq\eta|q(x_{j})|^{\alpha},~{}~{}\forall j\in[m].$ (2.4)
Unlike the original min-max problem (1.1), (2.4) is a standard minimization. The following theorem shows that, in general, the two infimums of (1.1) and (2.4) are consistent, and if (1.1) admits a solution, it can be recovered from the linearization (2.4). It is a generalization of [43, Theorem 2.1].
Theorem 2.2.
Given $m\geq{n_{1}}+n_{2}+2$ distinct nodes ${\cal X}=\{x_{j}\}_{j=1}^{m}$ on $\Omega\subset\mathbb{C}$, let $\eta_{\alpha}$ be the infimun of (2.4). If (2.4) has a solution $(\eta_{\alpha},\hat{p},\hat{q})$ with $\hat{p}\not\equiv 0$ and $\hat{q}\not\equiv 0$, then $\eta_{\alpha}=(\eta_{\infty})^{\alpha}$. Consequently, whenever (1.1) has an irreducible solution $\xi^{*}=p^{*}/q^{*}$, then if (2.4) admits a solution $(\eta_{\alpha},\hat{p},\hat{q})$ with $\hat{p}\not\equiv 0$ and $\hat{q}\not\equiv 0$, then $(p^{*},q^{*})$ can be solved from (2.4).
Proof.
First, it is true that $(\eta_{\alpha})^{\frac{1}{\alpha}}\leq\eta_{\infty}$. In fact, for any irreducible sequence $\{\xi^{(k)}=p^{(k)}/q^{(k)}\}$ satisfying $\eta_{\infty}^{(k)}:=\|\boldsymbol{f}-\xi^{(k)}(\boldsymbol{x})\|_{\infty}\rightarrow\eta_{\infty}$, the triplet $(\eta_{\infty}^{(k)},p^{(k)},q^{(k)})$ is feasible for (2.4), and thus $(\eta_{\alpha})^{\frac{1}{\alpha}}\leq\eta_{\infty}^{(k)}\rightarrow\eta_{\infty}$.
To show $(\eta_{\alpha})^{\frac{1}{\alpha}}=\eta_{\infty}$, suppose by contradiction that $(\eta_{\alpha})^{\frac{1}{\alpha}}<\eta_{\infty}$. Let $(\eta_{\alpha},\hat{p},\hat{q})$ be a solution for (2.4) with $\hat{p}\not\equiv 0$. Suppose without loss of generality that $\hat{q}(x_{j})=0$ for $j=1,\dots,s$. As $\hat{q}\not\equiv 0$, we have $s\leq n_{2}.$ The constraint of (2.4) leads to
$|f_{j}\hat{q}(x_{j})-\hat{p}(x_{j})|^{\alpha}=|\hat{p}(x_{j})|^{\alpha}\leq 0\Longrightarrow\hat{p}(x_{j})=0,\quad\forall j\in[s],$
implying $s\leq n_{1}$, because otherwise $\hat{p}\equiv 0$. This shows $s\leq\min(n_{1},n_{2})<m$. The proof is completed if $s=0$ (i.e., $q(x_{j})\neq 0$ for any $j\in[m]$) because
$|f(x_{j})\hat{q}(x_{j})-\hat{p}(x_{j})|^{\alpha}\leq\eta_{\alpha}|\hat{q}(x_{j})|^{\alpha}\Longrightarrow\left|f(x_{j})-\frac{\hat{p}(x_{j})}{\hat{q}(x_{j})}\right|\leq(\eta_{\alpha})^{\frac{1}{\alpha}}<\eta_{\infty},$
contradicting with the fact that $\eta_{\infty}$ is the infimum of (1.1). We next consider $s\geq 1$.
Choose a parameterized rational polynomial
$\frac{\hat{p}(x;\delta)}{\hat{q}(x;\delta)}=\frac{\hat{p}(x)+\delta p(x)}{\hat{q}(x)+\delta}\in\mathscr{R}_{(n_{1},n_{2})},~{}~{}\delta\in\mathbb{C},$
for some polynomial $p\in\mathbb{P}_{n_{1}}$. We will finish the proof by showing that there is a polynomial $p$ and sufficiently small $\delta$ so that $\hat{q}(x_{j};\delta)\neq 0$ for any $j\in[m]$ and
$\left\|f(\boldsymbol{x})-\frac{\hat{p}(\boldsymbol{x};\delta)}{\hat{q}(\boldsymbol{x};\delta)}\right\|_{\infty}<\eta_{\infty},$ (2.5)
which contradicts with the fact that $\eta_{\infty}$ is the infimum of (1.1).
To this end, we choose any polynomial $p\in\mathbb{P}_{n_{1}}$ so that $p(x_{j})=f(x_{j})$ for $j\in[s]$. Such a polynomial exists as $s\leq\min(n_{1},n_{2})$. Now, for any sufficiently small $\delta\neq 0$, we know that $\hat{q}(x_{j};\delta)\neq 0$ for any $j\in[m]$. Moreover, for any node $x_{j}$ with $\hat{q}(x_{j})\neq 0$ (i.e., $j\not\in[s]$), similarly, by $(\eta_{\alpha})^{\frac{1}{\alpha}}<\eta_{\infty}$, we have
$|f(x_{j})\hat{q}(x_{j})-\hat{p}(x_{j})|^{\alpha}\leq\eta_{\alpha}|\hat{q}(x_{j})|^{\alpha}\Longrightarrow\left|f(x_{j})-\frac{\hat{p}(x_{j})}{\hat{q}(x_{j})}\right|\leq(\eta_{\alpha})^{\frac{1}{\alpha}}<\eta_{\infty},$
and
$\left|\frac{\hat{p}(x_{j})+\delta p(x_{j})}{\hat{q}(x_{j})+\delta}-\frac{\hat{p}(x_{j})}{\hat{q}(x_{j})}\right|=\left|\frac{\delta p(x_{j})\hat{q}(x_{j})-\delta\hat{p}(x_{j})}{(\hat{q}(x_{j})+\delta)\hat{q}(x_{j})}\right|\rightarrow 0,~{}~{}{\rm as}~{}~{}\delta\rightarrow 0.$
Hence, for any sufficiently small $\delta$, it follows
$\left|\frac{\hat{p}(x_{j})+\delta p(x_{j})}{\hat{q}(x_{j})+\delta}-f(x_{j})\right|$ $\leq\left|\frac{\hat{p}(x_{j})+\delta p(x_{j})}{\hat{q}(x_{j})+\delta}-\frac{\hat{p}(x_{j})}{\hat{q}(x_{j})}\right|+\left|f(x_{j})-\frac{\hat{p}(x_{j})}{\hat{q}(x_{j})}\right|<\eta_{\infty}.$
On the other hand, for any node $x_{j}$ with $\hat{q}(x_{j})=0$ (i.e., $j\in[s]$), it holds from the constraint of (2.4) that $\hat{p}(x_{j})=0$ and
$\left|\frac{\hat{p}(x_{j})+\delta p(x_{j})}{\hat{q}(x_{j})+\delta}-f(x_{j})\right|=\left|p(x_{j})-f(x_{j})\right|=0\leq\eta_{\infty}.$
This leads to (2.5). Consequently, whenever (1.1) admits an irreducible solution $\xi^{*}=p^{*}/q^{*}$, and $(\eta_{\alpha},\hat{p},\hat{q})$ with $\hat{p}\not\equiv 0$ is a solution to (2.4), then we know that $(\eta_{\infty})^{\alpha}=\eta_{\alpha}$; also, the irreducible solution $\xi^{*}=p^{*}/q^{*}$ implies $|f_{j}q^{*}(x_{j})-p^{*}(x_{j})|^{\alpha}\leq\eta_{\alpha}|q^{*}(x_{j})|^{\alpha},~{}\forall j\in[m]$, showing that the triplet $((\eta_{\infty})^{\alpha},p^{*},q^{*})$ is also a solution of (2.4). ∎
2.2 A dual programming
Even though we have transformed the original bi-level min-max problem (1.1) into a standard optimization (as a primal programming), directly handling (2.4) is still hard. The idea in [43] is to develop the dual programming of (2.4), and then employ Lawson’s idea for the linear Chebyshev approximation for solving the dual programming.
For the computational convenience, we choose specially $\alpha=2$. Thus the dual function [43] of (2.4) can be given by
$d_{2}(\boldsymbol{w})$ $=\min_{\begin{subarray}{c}p\in\mathbb{P}_{n_{1}},~{}q\in\mathbb{P}_{n_{2}}\\ \sum_{j=1}^{m}w_{j}|q(x_{j})|^{2}=1\end{subarray}}\sum_{j=1}^{m}w_{j}|f_{j}q(x_{j})-p(x_{j})|^{2}$
$=\min_{\begin{subarray}{c}\boldsymbol{a}\in\mathbb{C}^{n_{1}+1},~{}\boldsymbol{b}\in\mathbb{C}^{n_{2}+1}\\ \|\sqrt{W}\Phi\boldsymbol{b}\|_{2}=1\end{subarray}}\left\|\sqrt{W}[-\Psi,F\Phi]\left[\begin{array}[]{c}\boldsymbol{a}\\ \boldsymbol{b}\end{array}\right]\right\|_{2}^{2},$ (2.8)
where $W={\rm diag}(\boldsymbol{w}),~{}F={\rm diag}(\boldsymbol{f})$. The following weak duality [43] has been proved:
$\forall\boldsymbol{w}\in{\cal S},~{}d_{2}(\boldsymbol{w})\leq(\eta_{\infty})^{2}\Longrightarrow\max_{\boldsymbol{w}\in{\cal S}}d_{2}(\boldsymbol{w})\leq(\eta_{\infty})^{2}.$ (2.9)
Moreover, by relying on Ruttan’s sufficient condition ([29, Theorems 2.1]; see also [19, Theorem 2] and [32, Theorem 3]), we can also have the following theoretical guarantee for solving the minimax approximant $\xi^{*}$.
Theorem 2.3 ([43]).
Given $m\geq{n_{1}}+n_{2}+2$ distinct nodes ${\cal X}=\{x_{j}\}_{j=1}^{m}$ on $\Omega$, we have the weak duality (2.9). Let $\boldsymbol{w}^{*}\in{\cal S}$ be the solution to the dual problem
$\max_{\boldsymbol{w}\in{\cal S}}d_{2}(\boldsymbol{w}),$ (2.10)
and $(\boldsymbol{a}^{*},\boldsymbol{b}^{*})$ achieve the minimum $d_{2}(\boldsymbol{w}^{*})$ of (2.8). Then if the associated rational polynomial $\xi^{*}(x)=\frac{p^{*}(x)}{q^{*}(x)}=\frac{[\psi_{0}(x),\dots,\psi_{n_{1}}(x)]\boldsymbol{a}^{*}}{[\phi_{0}(x),\dots,\phi_{n_{2}}(x)]\boldsymbol{b}^{*}}$ is irreducible and
$\|\boldsymbol{f}-\xi^{*}(\boldsymbol{x})\|_{\infty}=\sqrt{d_{2}(\boldsymbol{w}^{*})},$ (2.11)
then $\xi^{*}$ is the minimax approximant of (1.1). When (2.11) holds, then we also have the following complementary slackness property:
$w_{j}^{*}\left(\|\boldsymbol{f}-\xi^{*}(\boldsymbol{x})\|_{\infty}-|f_{j}-\xi^{*}(x_{j})|\right)=0,\quad\forall j=1,2,\dots,m.$ (2.12)
We remark that the condition (2.11) implies the strong duality $\max_{\boldsymbol{w}\in{\cal S}}d_{2}(\boldsymbol{w})=(\eta_{\infty})^{2}$ [43, Theorem 4.3], which is equivalent to Ruttan’s sufficient condition for the original rational minimax problem (1.1). Therefore, under Ruttan’s sufficient condition, the complementary slackness property (2.12) necessarily holds at the maximizer $\boldsymbol{w}^{*}$ of (2.8). Furthermore, in the framework of dual programming (2.8), the accuracy of the associated approximation $\xi$ corresponding to the minimization (2.8) at an approximation $\boldsymbol{w}$ of $\boldsymbol{w}^{*}$, can be measured by the relative error
$\epsilon(\boldsymbol{w}):=\left|\frac{\sqrt{d_{2}(\boldsymbol{w})}-e(\xi)}{e(\xi)}\right|.$
This serves as a stopping rule for Lawson’s iteration (Algorithm 1).
2.3 Optimality for the dual objective function
To compute the dual function $d_{2}(\boldsymbol{w})$, a minimization problem (2.8) needs to be solved. The following proposition provides the optimality condition for this minimization.
Proposition 2.1 ([43]).
For $\boldsymbol{w}\in{\cal S}$, we have
  • (i)
    $\boldsymbol{c}(\boldsymbol{w})=\left[\begin{array}[]{c}\boldsymbol{a}(\boldsymbol{w})\\ \boldsymbol{b}(\boldsymbol{w})\end{array}\right]\in\mathbb{C}^{n_{1}+n_{2}+2}$ is a solution of (2.10) if and only if it an eigenvector of the Hermitian positive semi-definite generalized eigenvalue problem $(A_{\boldsymbol{w}},B_{\boldsymbol{w}})$ and $d_{2}(\boldsymbol{w})$ is the smallest eigenvalue satisfying
    $A_{\boldsymbol{w}}\boldsymbol{c}(\boldsymbol{w})=d_{2}(\boldsymbol{w})B_{\boldsymbol{w}}\boldsymbol{c}(\boldsymbol{w})~{}\mbox{ and }~{}\boldsymbol{c}(\boldsymbol{w})^{\operatorname{H}}B_{\boldsymbol{w}}\boldsymbol{c}(\boldsymbol{w})=1,$ (2.13)
    where
    $A_{\boldsymbol{w}}:$ $=[-\Psi,F\Phi]^{\operatorname{H}}W[-\Psi,F\Phi]=\left[\begin{array}[]{cc}\Psi^{\operatorname{H}}W\Psi&-\Psi^{\operatorname{H}}FW\Phi\\ -\Phi^{\operatorname{H}}WF^{\operatorname{H}}\Psi&\Phi^{\operatorname{H}}F^{\operatorname{H}}WF\Phi\end{array}\right],$
    $B_{\boldsymbol{w}}:$ $=[0,\Phi]^{\operatorname{H}}W[0,\Phi]=\left[\begin{array}[]{cc}0&0\\ 0&\Phi^{\operatorname{H}}W\Phi\end{array}\right];$
  • (ii)
    the Hermitian matrix $H_{\boldsymbol{w}}:=A_{\boldsymbol{w}}-d_{2}(\boldsymbol{w})B_{\boldsymbol{w}}\succeq 0$, i.e., $H_{\boldsymbol{w}}$ is positive semi-definite;
  • (iii)
    let $W^{\frac{1}{2}}\Phi=Q_{q}R_{q}$ and $W^{\frac{1}{2}}\Psi=Q_{p}R_{p}$ be the thin QR factorizations where $Q_{q}\in\mathbb{C}^{m\times\widetilde{n}_{2}}$, $Q_{p}\in\mathbb{C}^{m\times\widetilde{n}_{1}}$, $R_{q}\in\mathbb{C}^{\widetilde{n}_{2}\times(n_{2}+1)}$, $R_{p}\in\mathbb{C}^{\widetilde{n}_{1}\times(n_{1}+1)}$ with $\widetilde{n}_{1}=\operatorname{rank}(W^{\frac{1}{2}}\Psi)$ and $\widetilde{n}_{2}=\operatorname{rank}(W^{\frac{1}{2}}\Phi)$. Then $(d_{2}(\boldsymbol{w}),R_{q}\boldsymbol{b}(\boldsymbol{w}))$ is an eigenpair associated with the smallest eigenvalue of the Hermitian positive semi-definite matrix $S_{F}-S_{qp}S_{qp}^{\operatorname{H}}\in\mathbb{C}^{\widetilde{n}_{2}\times\widetilde{n}_{2}}$ satisfying
    $(S_{F}-S_{qp}S_{qp}^{\operatorname{H}})R_{q}\boldsymbol{b}(\boldsymbol{w})=d_{2}(\boldsymbol{w})R_{q}\boldsymbol{b}(\boldsymbol{w}),~{}~{}\|R_{q}\boldsymbol{b}(\boldsymbol{w})\|_{2}=1,~{}~{}R_{p}\boldsymbol{a}(\boldsymbol{w})=S_{qp}^{\operatorname{H}}R_{q}\boldsymbol{b}(\boldsymbol{w}),$ (2.14)
    where
    $S_{F}=Q_{q}^{\operatorname{H}}|F|^{2}Q_{q}\in\mathbb{C}^{\widetilde{n}_{2}\times\widetilde{n}_{2}},~{}~{}S_{qp}=Q_{q}^{\operatorname{H}}F^{\operatorname{H}}Q_{p}\in\mathbb{C}^{\widetilde{n}_{2}\times\widetilde{n}_{1}}.$
    Moreover, assuming that $[Q_{p},Q_{p}^{\perp}]\in\mathbb{C}^{m\times m}$ is unitary, then $(\sqrt{d_{2}(\boldsymbol{w})},R_{q}\boldsymbol{b}(\boldsymbol{w}))$ is the right singular pair associated with the smallest singular value of both $(Q_{p}^{\perp})^{\operatorname{H}}FQ_{q}\in\mathbb{C}^{(m-\widetilde{n}_{1}-1)\times(\widetilde{n}_{2}+1)}$ and $(I_{m}-Q_{p}Q_{p}^{\operatorname{H}})FQ_{q}\in\mathbb{C}^{m\times(\widetilde{n}_{2}+1)}$.
As remarked in [43] that the normalized condition $\boldsymbol{c}(\boldsymbol{w})^{\operatorname{H}}B_{\boldsymbol{w}}\boldsymbol{c}(\boldsymbol{w})=1$ in (2.13) can always be fulfilled whenever $q\not\equiv 0$ and there are at least $n_{2}+1$ positive elements in $w_{j}$. In fact,
$0=\boldsymbol{c}(\boldsymbol{w})^{\operatorname{H}}B_{\boldsymbol{w}}\boldsymbol{c}(\boldsymbol{w})=\sum_{j=1}^{m}w_{j}|q(x_{j})|^{2}\Longrightarrow w_{j}q(x_{j})=0,~{}\forall j\in[m];$
thus, any node $x_{j}$ with $w_{j}>0$ is a zero of $q$, and if $\boldsymbol{w}$ has at least $n_{2}+1$ positive elements, it implies that $q\equiv 0$. Thus, in the following discussion, we assume that any $\boldsymbol{w}^{(k)}$ during iteration has at least $n_{2}+1$ positive elements.
We remark that, by rewriting the first $n_{1}+1$ rows and the last $n_{2}+1$ rows of the optimality condition $A_{\boldsymbol{w}}\boldsymbol{c}(\boldsymbol{w})=d_{2}(\boldsymbol{w})B_{\boldsymbol{w}}\boldsymbol{c}(\boldsymbol{w})$ in (2.13), we have
Corollary 2.1.
Let $\boldsymbol{p}=\Psi\boldsymbol{a}(\boldsymbol{w})$ and $\boldsymbol{q}=\Phi\boldsymbol{b}(\boldsymbol{w})$ be from the solution of (2.8) with the weight vector $\boldsymbol{w}$. Then
$F\boldsymbol{q}-\boldsymbol{p}\perp_{\boldsymbol{w}}{\rm span}(\Psi),~{}~{}F^{\operatorname{H}}(F\boldsymbol{q}-\boldsymbol{p})-d_{2}(\boldsymbol{w})\boldsymbol{q}\perp_{\boldsymbol{w}}{\rm span}(\Phi).$ (2.15)
Besides the optimality in solving the minimization for $d_{2}(\boldsymbol{w})$, we can further have the gradient of $d_{2}(\boldsymbol{w})$.
Proposition 2.2 ([43]).
For $\boldsymbol{w}>0$, let $d_{2}(\boldsymbol{w})$ be the smallest eigenvalue of the Hermitian positive semi-definite generalized eigenvalue problem (2.13), and $\boldsymbol{c}(\boldsymbol{w})=\left[\begin{array}[]{c}\boldsymbol{a}(\boldsymbol{w})\\ \boldsymbol{b}(\boldsymbol{w})\end{array}\right]\in\mathbb{C}^{n_{1}+n_{2}+2}$ be the associated eigenvector. Denote $\boldsymbol{p}=[p_{1},\dots,p_{m}]^{\operatorname{T}}=\Psi\boldsymbol{a}(\boldsymbol{w})\in\mathbb{C}^{m}$ and $\boldsymbol{q}=[q_{1},\dots,q_{m}]^{\operatorname{T}}=\Phi\boldsymbol{b}(\boldsymbol{w})\in\mathbb{C}^{m}$. If $d_{2}(\boldsymbol{w})$ is a simple eigenvalue, then $d_{2}(\boldsymbol{w})$ is differentiable with respect to $\boldsymbol{w}$ and its gradient is
$\nabla d_{2}(\boldsymbol{w})=\left[\begin{array}[]{c}|f_{1}q_{1}-p_{1}|^{2}-d_{2}(\boldsymbol{w})|q_{1}|^{2}\\ |f_{2}q_{2}-p_{2}|^{2}-d_{2}(\boldsymbol{w})|q_{2}|^{2}\\ \vdots\\ |f_{m}q_{m}-p_{m}|^{2}-d_{2}(\boldsymbol{w})|q_{m}|^{2}\end{array}\right]=:|F\boldsymbol{q}-\boldsymbol{p}|^{2}-d_{2}(\boldsymbol{w})|\boldsymbol{q}|^{2}\in\mathbb{R}^{m}.$ (2.16)
2.4 Lawson’s iteration
Within the framework of the dual programming, it has been claimed [43] that Lawson’s iteration is a method for solving the dual problem (2.10). For the rational minimax approximation, Lawson’s iteration [43] is implemented as in Algorithm 1. Numerical results have been reported in [43], indicating that Lawson’s iteration generally converges monotonically with respect to the dual function value $d_{2}(\boldsymbol{w})$.
Algorithm 1 A rational Lawson’s iteration [43] for (1.2)
0:  Given samples $\{(x_{j},f_{j})\}_{j=1}^{m}$ and $0\leq n_{1}+n_{2}+2\leq m$ with $x_{j}\in\Omega$, a relative tolerance for the strong duality $\epsilon_{r}>0$, the maximum number $k_{\rm maxit}$ of iterations;
1:  (Initialization) Let $k=0$; choose $0<\boldsymbol{w}^{(0)}\in{\cal S}$ and a tolerance $\epsilon_{w}$ for the weights;
2:  (Filtering) Remove nodes $x_{i}$ with $w_{i}^{(k)}<\epsilon_{w}$;
3:  Compute $d_{2}(\boldsymbol{w}^{(k)})$ and the associated vector $\xi^{(k)}(\boldsymbol{x})=p^{(k)}(\boldsymbol{x})./q^{(k)}(\boldsymbol{x})$ according to Proposition 2.1;
4:  (Stop rule) Stop either if $k\geq k_{\rm maxit}$ or
$\epsilon(\boldsymbol{w}^{(k)}):=\left|\frac{\sqrt{d_{2}(\boldsymbol{w}^{(k)})}-e(\xi^{{(k)}})}{e(\xi^{{(k)}})}\right|<\epsilon_{r},~{}~{}{\rm where}~{}~{}e(\xi^{{(k)}})=\|\boldsymbol{f}-\xi^{(k)}(\boldsymbol{x})\|_{\infty};$
5:  (Updating weights) Update the weight vector $\boldsymbol{w}^{(k+1)}$ according to
$w_{j}^{(k+1)}=\frac{w_{j}^{(k)}\left|f_{j}-\xi^{(k)}(x_{j})\right|^{\beta}}{\sum_{i}w_{i}^{(k)}\left|f_{i}-\xi^{(k)}(x_{j})\right|^{\beta}},~{}~{}\forall j,$ (2.17)
with the Lawson exponent $\beta>0$, and goto Step 2 with $k=k+1$.
Remark 2.1.
It is interesting to point out that Lawson’s iteration in Algorithm 1 naturally reduces to the classical Lawson’s iteration [21] for the linear (polynomial) minimax approximation when $n_{2}=0$, which corresponds to $q\equiv 1$ and $\boldsymbol{q}^{\operatorname{H}}W\boldsymbol{q}\equiv 1$. This observation unifies our following convergence analysis.
In practice, the implementation of Step 3 should handle the stability and accuracy for computing $d_{2}(\boldsymbol{w}^{(k)})$ and the associated vector $\xi^{(k)}(\boldsymbol{x})=p^{(k)}(\boldsymbol{x})./q^{(k)}(\boldsymbol{x})$. In [43], the Vandermonde with Arnoldi process [4, 42, 18] is employed for this step.
It is noticed that Lawson’s iteration (2.17) relies on the error vector $\boldsymbol{f}-\xi^{(k)}(\boldsymbol{x})$ to update the weights. In this procedure, each entry of the denominator vector $\boldsymbol{q}^{(k)}$ is assumed to be nonzero; this is generally the case in practice. Even if in the extreme situation where some of entries of $\boldsymbol{q}^{(k)}$ vanish, we can find a perturbation of $\boldsymbol{q}^{(k)}$ using the right singular vectors of $(Q_{p}^{\perp})^{\operatorname{H}}FQ_{q}\in\mathbb{C}^{(m-\widetilde{n}_{1}-1)\times(\widetilde{n}_{2}+1)}$ or $(I-Q_{p}Q_{p}^{\operatorname{H}})FQ_{q}\in\mathbb{C}^{m\times(\widetilde{n}_{2}+1)}$ corresponding to the smallest and second smallest singular values (refer to (iii) of Proposition 2.1). Therefore, in our following discussion, we assume the following two assumptions:
  • (A1)
    $q_{j}^{(k)}=q^{(k)}(x_{j})\neq 0$ for any $j\in[m]$ and any $k\geq 0$;
  • (A2)
    the cardinality $|{\cal I}_{\boldsymbol{w}^{(k)}}|\geq\max\{n_{1}+1,n_{2}+1\}$, where ${\cal I}_{\boldsymbol{w}^{(k)}}:=\{j\in[m]|w_{j}^{(k)}\neq 0\}$.
Proposition 2.3.
If $d_{2}(\boldsymbol{w}^{(k)})>0$, then
${\cal I}_{\boldsymbol{w}^{(k+1)}}=\{j\in[m]|w_{j}^{(k)}r_{j}^{(k)}\neq 0\},~{}~{}{\rm where}~{}~{}r_{j}^{(k)}=|f(x_{j})-\xi^{(k)}(x_{j})|.$
Proof.
By (A1), the condition $d_{2}(\boldsymbol{w}^{(k)})>0$ implies $\sum_{i}w_{i}^{(k)}\left|f_{i}-\xi^{(k)}(x_{j})\right|^{\beta}>0$. Thus, according to Lawson’s updating rule (2.17), the conclusion follows due to the fact that $w_{j}^{(k+1)}=0$ is equivalent to $w_{j}^{(k)}r_{j}^{(k)}=0$. ∎
3 A lower bound of the dual objective function value
We first relate Lawson’s iteration with the gradient ascent direction. Indeed, when the current $0<\boldsymbol{w}\in{\cal S}$, $q_{j}=\boldsymbol{q}^{\operatorname{T}}\boldsymbol{e}_{j}\neq 0~{}\forall j\in[m]$ and $d(\boldsymbol{w})$ is the simple eigenvalue of $(A_{\boldsymbol{w}},B_{\boldsymbol{w}})$, then we know that the direction
$\boldsymbol{g}(\boldsymbol{w}):={\rm diag}\left(\frac{w_{1}}{|q_{1}|^{2}},\ldots,\frac{w_{m}}{|q_{m}|^{2}}\right)\nabla d_{2}(\boldsymbol{w})\in\mathbb{R}^{m}$
is an ascent direction for the dual function $d_{2}(\boldsymbol{w})$ because by (2.16)
$\boldsymbol{g}(\boldsymbol{w})^{\operatorname{T}}\nabla d_{2}(\boldsymbol{w})=\nabla d_{2}(\boldsymbol{w})^{\operatorname{T}}{\rm diag}\left(\frac{w_{1}}{|q_{1}|^{2}},\ldots,\frac{w_{m}}{|q_{m}|^{2}}\right)\nabla d_{2}(\boldsymbol{w})>0.$
Note that using this direction, the updating with the step-size $s=\frac{1}{d_{2}(\boldsymbol{w})}>0$ gives
$\widetilde{\boldsymbol{w}}:=\boldsymbol{w}+s\boldsymbol{g}={\rm diag}(\boldsymbol{w})|\boldsymbol{f}-\xi(\boldsymbol{x})|^{2}/d_{2}(\boldsymbol{w})$
which, after scaling it to have $\widetilde{\boldsymbol{w}}\in{\cal S}$, implies the iteration is the same as Lawson’s iteration (2.17) with $\beta=2$. Note that the scaling $\widetilde{\boldsymbol{w}}\leftarrow\frac{\widetilde{\boldsymbol{w}}}{\boldsymbol{e}^{\operatorname{T}}\widetilde{\boldsymbol{w}}}$ can be viewed as a certain projection onto ${\cal S}$. From this point of view, we can say that a Lawson’s iteration in Algorithm 1 with the Lawson exponent $\beta=2$ is just an ascent gradient step with a specific step-size followed by a certain projection onto ${\cal S}$.
To see the more clear relation between the two successive objective values $d_{2}(\boldsymbol{w}^{(k)})$ and $d_{2}(\boldsymbol{w}^{(k+1)})$, for $\boldsymbol{w}\geq 0$, we introduce the $\boldsymbol{w}$-inner (positive semidefinite) product defined by $\langle\boldsymbol{y},\boldsymbol{z}\rangle_{\boldsymbol{w}}=\boldsymbol{y}^{\operatorname{H}}W\boldsymbol{z}$ and $\|\boldsymbol{y}\|_{\boldsymbol{w}}=\sqrt{\boldsymbol{y}^{\operatorname{H}}W\boldsymbol{y}}$, where $W={\rm diag}(\boldsymbol{w})$. The following lemma, which is a generalized result of the standard least-squares problem, plays an important role in establishing the convergence of Lawson’s iteration [21] for the linear minimax approximation problem [26, Lemma 13-12].
Lemma 3.1.
Given $0\leq\boldsymbol{w}\in\mathbb{R}^{m}$ satisfying $|{\cal I}_{\boldsymbol{w}}|\geq n$ with ${\cal I}_{\boldsymbol{w}}=\{j|w_{j}>0\}$, a matrix $A\in\mathbb{C}^{m\times n}$ with $\operatorname{rank}(A({\cal I}_{\boldsymbol{w}},:))=n$ and $\boldsymbol{z}\in\mathbb{C}^{m}$, let $\boldsymbol{x}_{*}$ be the solution to the least-squares problem
$\min_{\boldsymbol{x}\in{\rm span}(A)}\|\boldsymbol{x}-\boldsymbol{z}\|_{\boldsymbol{w}}.$
Then we have
$\frac{\boldsymbol{x}_{*}-\boldsymbol{z}}{\|\boldsymbol{x}_{*}-\boldsymbol{z}\|_{\boldsymbol{w}}}\in\arg\max_{\|\boldsymbol{y}\|_{\boldsymbol{w}}=1,~{}\boldsymbol{y}^{\operatorname{H}}WA=0}|\langle\boldsymbol{z},\boldsymbol{y}\rangle_{\boldsymbol{w}}|.$
Proof.
We first consider the case $\boldsymbol{w}>0$.
Denote by $P\in\mathbb{C}^{m\times(m-n)}$ the $\boldsymbol{w}$-orthogonal basis for the complement [20, Chapter 1.6.3] of ${\rm span}(A)$ satisfying ${\rm span}(P)\oplus{\rm span}(A)=\mathbb{C}^{m}$, $P^{\operatorname{H}}WP=I_{m-n}$ and $P^{\operatorname{H}}WA=0$. Note that $PP^{\operatorname{H}}W$ and $I-PP^{\operatorname{H}}W$ are projections onto ${\rm span}(P)$ and ${\rm span}(A)$, respectively. For the least-squares problem, we know that $\boldsymbol{x}_{*}$ is the solution if and only if
$(\boldsymbol{x}_{*}-\boldsymbol{z})^{\operatorname{H}}WA=0,~{}~{}(i.e.,~{}\boldsymbol{x}_{*}-\boldsymbol{z}=PP^{\operatorname{H}}W\boldsymbol{z}).$
Note $\|\boldsymbol{x}_{*}-\boldsymbol{z}\|_{\boldsymbol{w}}=\|PP^{\operatorname{H}}W\boldsymbol{z}\|_{\boldsymbol{w}}=\|P^{\operatorname{H}}W\boldsymbol{z}\|_{2}$. Also, the constraint $\|\boldsymbol{y}\|_{\boldsymbol{w}}=1,~{}\boldsymbol{y}^{\operatorname{H}}WA=0$ can be parameterized as $\boldsymbol{y}=P\boldsymbol{t}$ for $\boldsymbol{t}\in\mathbb{C}^{m-n}$ with $\|\boldsymbol{t}\|_{2}=1$. Since
$|\langle\boldsymbol{z},\boldsymbol{y}\rangle_{\boldsymbol{w}}|=|\boldsymbol{z}^{\operatorname{H}}WP\boldsymbol{t}|\leq\|P^{\operatorname{H}}W\boldsymbol{z}\|_{2},$
and the equality holds if $\boldsymbol{t}=P^{\operatorname{H}}W\boldsymbol{z}/\|P^{\operatorname{H}}W\boldsymbol{z}\|_{2}$, it implies that
$\boldsymbol{y}=\frac{PP^{\operatorname{H}}W\boldsymbol{z}}{\|P^{\operatorname{H}}W\boldsymbol{z}\|_{2}}=\frac{\boldsymbol{x}_{*}-\boldsymbol{z}}{\|\boldsymbol{x}_{*}-\boldsymbol{z}\|_{\boldsymbol{w}}}$
is the maximizer and the conclusion follows.
For the general case $\boldsymbol{w}\geq 0$, assume ${\cal I}_{\boldsymbol{w}}=\{1,2,\dots,t\}=[t]$, and partition accordingly $\boldsymbol{x}=[\boldsymbol{x}_{1}^{\operatorname{T}},\boldsymbol{x}_{2}^{\operatorname{T}}]^{\operatorname{T}}$, $\boldsymbol{z}=[\boldsymbol{z}_{1}^{\operatorname{T}},\boldsymbol{z}_{2}^{\operatorname{T}}]^{\operatorname{T}}$, $\boldsymbol{w}=[\boldsymbol{w}_{1}^{\operatorname{T}},\boldsymbol{w}_{2}^{\operatorname{T}}]^{\operatorname{T}}$, $\boldsymbol{y}=[\boldsymbol{y}_{1}^{\operatorname{T}},\boldsymbol{y}_{2}^{\operatorname{T}}]^{\operatorname{T}}$, and $A=[A_{1}^{\operatorname{T}},A_{2}^{\operatorname{T}}]^{\operatorname{T}}$ with $A_{1}\in\mathbb{C}^{t\times n}$ and $\operatorname{rank}(A_{1})=n$. Note
$\min_{\boldsymbol{x}\in{\rm span}(A)}\|\boldsymbol{x}-\boldsymbol{z}\|_{\boldsymbol{w}}=\min_{\boldsymbol{x}_{1}\in{\rm span}(A_{1})}\|\boldsymbol{x}_{1}-\boldsymbol{z}_{1}\|_{\boldsymbol{w}_{1}},$
and the solution $\boldsymbol{x}^{*}$ is unique by assumptions. Indeed, $\boldsymbol{x}^{*}$ satisfies $(\boldsymbol{x}_{*}-\boldsymbol{z})^{\operatorname{H}}WA=0.$
On the other hand,
$\max_{\|\boldsymbol{y}\|_{\boldsymbol{w}}=1,~{}\boldsymbol{y}^{\operatorname{H}}WA=0}|\langle\boldsymbol{z},\boldsymbol{y}\rangle_{\boldsymbol{w}}|=\max_{\|\boldsymbol{y}_{1}\|_{\boldsymbol{w}_{1}}=1,~{}\boldsymbol{y}_{1}^{\operatorname{H}}W_{1}A_{1}=0}|\langle\boldsymbol{z}_{1},\boldsymbol{y}_{1}\rangle_{\boldsymbol{w}_{1}}|$
and by the proof of the first case $\boldsymbol{w}>0$, we know that any maximizer $\boldsymbol{y}=[\boldsymbol{y}_{1}^{\operatorname{T}},\boldsymbol{y}_{2}^{\operatorname{T}}]^{\operatorname{T}}$ takes $\boldsymbol{y}_{1}=\frac{\boldsymbol{x}{{}_{1}}_{*}-\boldsymbol{z}_{1}}{\|\boldsymbol{x}{{}_{1}}_{*}-\boldsymbol{z}_{1}\|_{\boldsymbol{w}_{1}}}$ and $\forall\boldsymbol{y}_{2}\in\mathbb{C}^{m-t}$. Thus, the conclusion follows. ∎
Lemma 3.2.
Let $(\boldsymbol{p}^{(k)}=\Psi\boldsymbol{a}^{(k)},~{}\boldsymbol{q}^{(k)}=\Phi\boldsymbol{b}^{(k)})$ be the solution for (2.8) with the weight $\boldsymbol{w}^{(k)}\in{\cal S}$ and $\xi^{(k)}(\boldsymbol{x})=\boldsymbol{p}^{(k)}./\boldsymbol{q}^{(k)}(\boldsymbol{x})$. If $d_{2}(\boldsymbol{w}^{(k)})>0$, then
$(F\boldsymbol{q}^{(k)}-\boldsymbol{p}^{(k)})./|\boldsymbol{f}-\xi^{(k)}(\boldsymbol{x})|^{{\beta}}\perp_{\boldsymbol{w}^{(k+1)}}{\rm span}(\Psi).$ (3.1)
Here we set $\frac{0}{0}=0$ for convenience.
Proof.
For simplicity, we denote the pairs $(\boldsymbol{p}^{(k)},\boldsymbol{q}^{(k)})$ and $(\boldsymbol{p}^{(k+1)},\boldsymbol{q}^{(k+1)})$ by $(\boldsymbol{p},\boldsymbol{q})$ and $(\widetilde{\boldsymbol{p}},\widetilde{\boldsymbol{q}})$ at the $k$th and $(k+1)$th step, respectively; this convenience applies to other quantities.
Notice that $d_{2}(\boldsymbol{w})>0$ implies
${\cal I}_{\widetilde{\boldsymbol{w}}}=\{j\in[m]|\widetilde{w}_{j}\neq 0\}=\{j\in[m]|w_{j}r_{j}\neq 0\}$
by Proposition 2.3. As $(\widetilde{\boldsymbol{a}},\widetilde{\boldsymbol{b}})$ is the solution to (2.8) associated with the weight $\widetilde{\boldsymbol{w}}$, for the given $\widetilde{\boldsymbol{b}}$, the vector $\widetilde{\boldsymbol{a}}$ is the solution to the following least-squares problem
$\min_{\boldsymbol{a}\in\mathbb{C}^{n_{1}+1}}\left\|F\widetilde{\boldsymbol{q}}-\Psi\boldsymbol{a}\right\|_{\widetilde{\boldsymbol{w}}}=\min_{\boldsymbol{z}\in{\rm span}(\Psi)}\left\|F\widetilde{\boldsymbol{q}}-\boldsymbol{z}\right\|_{\widetilde{\boldsymbol{w}}};$
thus we have the optimality $(F\widetilde{\boldsymbol{q}}-\widetilde{\boldsymbol{p}})^{\operatorname{H}}\widetilde{W}\Psi=0$ (see (2.15)); analogously, for the previous iteration, we have $(F\boldsymbol{q}-\boldsymbol{p})^{\operatorname{H}}W\Psi=0$ implying that $\forall\boldsymbol{t}\in{\rm span}(\Psi)$,
$0=\langle\boldsymbol{t},F\boldsymbol{q}-\boldsymbol{p}\rangle_{\boldsymbol{w}}$ $=\sum_{j=1}^{m}w_{j}\bar{t}_{j}q_{j}(f_{j}-\xi_{j})$
$=\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}\bar{t}_{j}q_{j}(f_{j}-\xi_{j})$
$=\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}\bar{t}_{j}q_{j}\operatorname{sign}(f_{j}-\xi_{j})|f_{j}-\xi_{j}|$
$=\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|f_{j}-\xi_{j}|^{\beta}\bar{t}_{j}q_{j}\operatorname{sign}(f_{j}-\xi_{j})|f_{j}-\xi_{j}|^{1-\beta}$
$=\gamma_{\beta}\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}\widetilde{w}_{j}\bar{t}_{j}q_{j}(f_{j}-\xi_{j})|f_{j}-\xi_{j}|^{-\beta}$
$=\gamma_{\beta}\sum_{j=1}^{m}\widetilde{w}_{j}\bar{t}_{j}q_{j}(f_{j}-\xi_{j})|f_{j}-\xi_{j}|^{-\beta}\quad\quad\mbox{(by ~{}$\widetilde{w}_{j}=0~{}\forall j\not\in{\cal I}_{\widetilde{\boldsymbol{w}}}$)}$
$=\gamma_{\beta}\langle\boldsymbol{t},(F\boldsymbol{q}-\boldsymbol{p})./|\boldsymbol{f}-\xi(\boldsymbol{x})|^{{\beta}}\rangle_{\widetilde{\boldsymbol{w}}};$
i.e., $(F\boldsymbol{q}-\boldsymbol{p})./|\boldsymbol{f}-\xi(\boldsymbol{x})|^{{\beta}}\perp_{\widetilde{\boldsymbol{w}}}{\rm span}(\Psi)$. ∎
We next provide a lower bound for the dual objective function value $d_{2}(\boldsymbol{w}^{(k+1)})$. This lower bound plays a crucial role in finding the near-optimal $\beta$ for Lawson’s iteration in the linear minimax approximation, and also is a key to establish monotonic convergence $d_{2}(\boldsymbol{w}^{(k+1)})\geq d_{2}(\boldsymbol{w}^{(k)})$ in the minimax rational approximation.
Theorem 3.1.
Let $(\boldsymbol{p}^{(k)}=\Psi\boldsymbol{a}^{(k)},\boldsymbol{q}^{(k)}=\Phi\boldsymbol{b}^{(k)})$ be the solution for (2.8) with the weight $\boldsymbol{w}^{(k)}\in{\cal S}$. Then
$\sqrt{d_{2}(\boldsymbol{w}^{(k+1)})}\geq d_{2}(\boldsymbol{w}^{(k)})\frac{\left|(\boldsymbol{q}^{(k+1)})^{\operatorname{H}}W^{(k)}\boldsymbol{q}^{(k)}\right|}{\gamma_{\beta}~{}\zeta_{\beta}},$ (3.2)
where $W^{(k)}={\rm diag}(\boldsymbol{w}^{(k)})$,
$\gamma_{\beta}=\left\||\boldsymbol{f}-\xi^{(k)}(\boldsymbol{x})|^{\frac{\beta}{2}}\right\|_{\boldsymbol{w}^{(k)}}^{2}:=\sum_{j=1}^{m}w_{j}^{(k)}\left|f_{j}-\xi^{(k)}(x_{j})\right|^{\beta},~{}~{}~{}~{}\zeta_{\beta}=\left\|\frac{(\boldsymbol{f}\boldsymbol{q}^{(k)}-\boldsymbol{p}^{(k)}).}{\left|\boldsymbol{f}-\xi^{(k)}(\boldsymbol{x})\right|^{{\beta}}}\right\|_{\boldsymbol{w}^{(k+1)}},$
and we set $\frac{0}{0}=0$ for convenience.
Proof.
Following the notation in the proof of Lemma 3.2, we omit the subscript $k$ in each quantity related with the $k$th iteration. The result of (3.2) is trivial if $d_{2}(\boldsymbol{w})=0$, and we assume $d_{2}(\boldsymbol{w})>0$, which, by Proposition 2.3, implies ${\cal I}_{\widetilde{\boldsymbol{w}}}=\{j\in[m]|\widetilde{w}_{j}\neq 0\}=\{j\in[m]|w_{j}r_{j}\neq 0\}.$
By assumptions (A1), (A2), $\widetilde{w}_{j}=w_{j}|f_{j}-\xi_{j}|^{\beta}/\gamma_{\beta}$, and the fact that any $n_{1}+1$ rows of $\Psi$ are linearly independent, we can apply Lemmas 3.1 and 3.2 to get
$d_{2}(\widetilde{\boldsymbol{w}})$ $=\langle F\widetilde{\boldsymbol{q}}-\widetilde{\boldsymbol{p}},F\widetilde{\boldsymbol{q}}-\widetilde{\boldsymbol{p}}\rangle_{\widetilde{\boldsymbol{w}}}$
$=\left|\left\langle F\widetilde{\boldsymbol{q}}-\widetilde{\boldsymbol{p}},F\widetilde{\boldsymbol{q}}\right\rangle_{\widetilde{\boldsymbol{w}}}\right|$
$=\sqrt{d_{2}(\widetilde{\boldsymbol{w}})}~{}\left|\left\langle\frac{F\widetilde{\boldsymbol{q}}-\widetilde{\boldsymbol{p}}}{\|F\widetilde{\boldsymbol{q}}-\widetilde{\boldsymbol{p}}\|_{\widetilde{\boldsymbol{w}}}},F\widetilde{\boldsymbol{q}}\right\rangle_{\widetilde{\boldsymbol{w}}}\right|$
$=\sqrt{d_{2}(\widetilde{\boldsymbol{w}})}~{}\max_{\|\boldsymbol{y}\|_{\widetilde{\boldsymbol{w}}}=1,~{}\boldsymbol{y}^{\operatorname{H}}\widetilde{W}\Psi=0}\left|\left\langle F\widetilde{\boldsymbol{q}},\boldsymbol{y}\right\rangle_{\widetilde{\boldsymbol{w}}}\right|$
[…]
$=\frac{\sqrt{d_{2}(\widetilde{\boldsymbol{w}})}}{\zeta_{\beta}}~{}\left|\sum_{j=1}^{m}\widetilde{w}_{j}\overline{f_{j}\widetilde{q_{j}}}(f_{j}q_{j}-p_{j})/|f_{j}-\xi_{j}|^{\beta}\right|$
$=\frac{\sqrt{d_{2}(\widetilde{\boldsymbol{w}})}}{\gamma_{\beta}\zeta_{\beta}}~{}\left|\sum_{j=1}^{m}w_{j}\overline{f_{j}\widetilde{q_{j}}}(f_{j}q_{j}-p_{j})\right|$
$=\frac{\sqrt{d_{2}(\widetilde{\boldsymbol{w}})}}{\gamma_{\beta}\zeta_{\beta}}~{}\left|\langle F\widetilde{\boldsymbol{q}},F\boldsymbol{q}-\boldsymbol{p}\rangle_{\boldsymbol{w}}\right|$
$=\frac{\sqrt{d_{2}(\widetilde{\boldsymbol{w}})}}{\gamma_{\beta}\zeta_{\beta}}{d_{2}(\boldsymbol{w})\left|\widetilde{\boldsymbol{q}}^{\operatorname{H}}W\boldsymbol{q}\right|},$
where the last equality $\langle F\widetilde{\boldsymbol{q}},F\boldsymbol{q}-\boldsymbol{p}\rangle_{\boldsymbol{w}}=d_{2}(\boldsymbol{w})\widetilde{\boldsymbol{q}}^{\operatorname{H}}W\boldsymbol{q}$ is due to the second optimality in (2.15) for the pair $(\boldsymbol{p},\boldsymbol{q})$:
$F^{\operatorname{H}}(F\boldsymbol{q}-\boldsymbol{p})-d(\boldsymbol{w})\boldsymbol{q}\perp_{\boldsymbol{w}}{\rm span}(\Phi)$ $\Longrightarrow\widetilde{\boldsymbol{b}}^{\operatorname{H}}\Phi^{\operatorname{H}}W(F^{\operatorname{H}}(F\boldsymbol{q}-\boldsymbol{p})-d_{2}(\boldsymbol{w})\boldsymbol{q})=0$
$\Longrightarrow\langle F\widetilde{\boldsymbol{q}},F\boldsymbol{q}-\boldsymbol{p}\rangle_{\boldsymbol{w}}=d_{2}(\boldsymbol{w})\widetilde{\boldsymbol{q}}^{\operatorname{H}}W\boldsymbol{q}.$
The proof is complete. ∎
4 $\beta=1$ is near-optimal for the linear minimax approximation
Our strategy for defining the optimal parameter $\beta$ at the $k$th iteration is based on the lower bound (3.2). We remark that $\widetilde{\boldsymbol{p}}$ and $\widetilde{\boldsymbol{q}}$ play different roles in Lawson’s iteration (Algorithm 1). In fact, recalling the dual function (2.8) or the optimality condition (2.15), we know that $\widetilde{\boldsymbol{p}}$ is essentially from a least-squares problem for the given $\widetilde{\boldsymbol{q}}$, which is essentially linearly dependent on the data $F,\widetilde{\boldsymbol{q}},\widetilde{\boldsymbol{w}}$; however, as $\widetilde{\boldsymbol{q}}$ both appear in the constraint and the objective function in (2.8), $\widetilde{\boldsymbol{q}}$ is related with an eigenvector (refer to (2.14)) of a matrix associated with the data $F,\widetilde{\boldsymbol{p}},\widetilde{\boldsymbol{w}}$, and therefore, is nonlinearly dependent on these data. For the lower bound (3.2), it is interesting to notice that only the numerator depends on the solution pair $(\widetilde{\boldsymbol{p}},\widetilde{\boldsymbol{q}})$ at new $\widetilde{\boldsymbol{w}}$, while the denominator $\gamma_{\beta}\zeta_{\beta}$ is only related with the information $\boldsymbol{p},\boldsymbol{q},\boldsymbol{w}$ at the current iteration. For the polynomial minimax approximation (i.e., $n_{2}=0$), particularly, as $\boldsymbol{q}^{(k+1)}\equiv\boldsymbol{q}^{(k)}$ and $(\boldsymbol{q}^{(k+1)})^{\operatorname{H}}W^{(k)}\boldsymbol{q}^{(k)}=1$ for all $k$, the lower bound in (3.2) only depends on $\boldsymbol{p}^{(k)},\boldsymbol{q}^{(k)},\boldsymbol{w}^{(k)}$. Based on this observation, we may define a near-optimal Lawson exponent $\beta_{*}^{(k)}$ as the minimizer of the lower bound (3.2), i.e.,
$\beta_{*}^{(k)}=\operatorname*{argmin}_{\beta\in\mathbb{R}}\zeta_{\beta}^{2}\gamma_{\beta}^{2}.$ (4.1)
Recalling $r_{j}=|f_{j}-\xi^{(k)}(x_{j})|$ and ${\cal I}_{\widetilde{\boldsymbol{w}}}=\{j\in[m]|\widetilde{w}_{j}>0\}=\{j\in[m]|w_{j}r_{j}>0\}$ in Proposition 2.3, we have
$\nu(\beta):=\zeta_{\beta}^{2}\gamma_{\beta}^{2}=\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}\right).$ (4.2)
Proposition 4.1.
The function $\nu(\beta)$ given in (4.2) is convex.
Proof.
First, note that the derivative of $\nu(\beta)$ is
$\nu^{\prime}(\beta)=(\zeta_{\beta}^{2}\gamma_{\beta}^{2})^{\prime}$ $=-\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}\log r_{j}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}\right)$
$+\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}\log r_{j}\right).$
Moreover,
$\nu^{\prime\prime}(\beta)=$ $\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}(\log r_{j})^{2}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}\right)+\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}(\log r_{j})^{2}\right)$
$-2\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}\log r_{j}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}\log r_{j}\right)$
$\geq$ $2\sqrt{\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}(\log r_{j})^{2}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}(\log r_{j})^{2}\right)}$
$-2\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}|\log r_{j}|\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}|\log r_{j}|\right)$
$\geq$ $2{\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}\left|q_{j}\right|^{2}r_{j}^{2-\beta}|\log r_{j}|\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}|\log r_{j}|\right)}$ (4.3)
$-2\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}|\log r_{j}|\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}|\log r_{j}|\right)$
$=$ $0,$
where the first inequality follows by using $a^{2}+b^{2}\geq 2\sqrt{|ab|}$, while the second is due to $\|\boldsymbol{a}\|_{2}\|\boldsymbol{b}\|_{2}\geq|\boldsymbol{a}^{\operatorname{T}}\boldsymbol{b}|$. This implies that $\nu(\beta)$ is convex. ∎
Proposition 4.2.
For the polynomial minimax approximation, we have
  • (i)
    $\beta=1$ is the global minimizer of (4.1). In this sense, $\beta=1$ achieves the maximum of the lower bound in (3.2) and can be viewed as the near-optimal Lawson exponent in Lawson’s iteration;
  • (ii)
    there is a $\beta_{0}>1$ so that for any $\beta\in(0,\beta_{0})$, the sequence of Lawson’s iteration satisfies $d_{2}(\boldsymbol{w}^{(k+1)})\geq d_{2}(\boldsymbol{w}^{(k)})$.
Proof.
For this special case of $\beta=1$ and $q\equiv 1$, we only need to notice
$\nu^{\prime}(1)$ $=-\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}\log r_{j}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}\right)+\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}\log r_{j}\right)=0,$
and thus the conclusion (i) follows from Proposition 4.1.
For (ii), by the convexity of $\nu(\beta)$ (i.e., $\nu^{\prime\prime}(\beta)\geq 0$) and ${{d_{2}(\boldsymbol{w})}}=\nu(0)\geq\nu(1)=\min_{\beta\in\mathbb{R}}\nu(\beta)$, we know that there is a $\beta_{0}>1$ so that $\nu(\beta_{0})=\nu(0)=d_{2}(\boldsymbol{w})$. Thus, for any $\beta\in(0,\beta_{0})$, it holds that $\nu(\beta)\leq{d_{2}(\boldsymbol{w})}$, which, according to the lower bound of $d_{2}(\widetilde{\boldsymbol{w}})$ in (3.2), leads to $d_{2}(\widetilde{\boldsymbol{w}})\geq d_{2}(\boldsymbol{w})$. ∎
Proposition 4.2 establishes the monotonic convergence and locally the near-optimal choice of the Lawson exponent at each iteration. For the global convergence and other convergence behaviors of Lawson’s iteration with $\beta=1$, the reader is referred to [21, 5, 27, 9, 26].
5 Monotonic convergence and complementary slackness for the rational minimax approximation
We now consider the convergence of Lawson’s iteration for the rational minimax approximation. Different from the linear case, a difficulty arises from the numerator $(\boldsymbol{q}^{(k+1)})^{\operatorname{H}}W^{(k)}\boldsymbol{q}^{(k)}$ of the lower bound (3.2), in which $\boldsymbol{q}^{(k+1)}$ is also dependent on the Lawson exponent $\beta$. Explicitly expressing $\boldsymbol{q}^{(k+1)}$ in terms of $\beta$ in general is impossible because $\boldsymbol{q}^{(k+1)}$ is related with an eigenvector of the matrix pencil $(A_{\boldsymbol{w}},B_{\boldsymbol{w}})$. However, locally around $\beta=0$, it is possible to analyze the term $(\boldsymbol{q}^{(k+1)})^{\operatorname{H}}W^{(k)}\boldsymbol{q}^{(k)}$, and therefore, the lower bound in (3.2) with respect to $\beta$. Based on this observation, we can conclude that, generically, for any sufficiently small $\beta$, the monotonic property $d_{2}(\boldsymbol{w}^{(k+1)})\geq d_{2}(\boldsymbol{w}^{(k)})$ holds. This convergence result is consistent with the numerical experiments of AAA-Lawson iteration where it is observed [11] that “taking $\beta$ to be smaller makes the algorithm much more robust”.
To develop our convergence, we need the following lemma.
Lemma 5.1.
Given $\boldsymbol{s}=[s_{1},\dots,s_{m}]^{\operatorname{T}}\in{\cal S}$, let $t(x)$ and $h(x)$ be strictly monotonically increasing on the interval $(a,b)$. Then for any $m$ points $x_{j}\in(a,b),~{}1\leq j\leq m$, we have
$\left(\sum_{j=1}^{m}s_{j}t(x_{j})\right)\left(\sum_{j=1}^{m}s_{j}h(x_{j})\right)\leq\sum_{j=1}^{m}s_{j}t(x_{j})h(x_{j});$ (5.1)
moreover, the equality in (5.1) holds if and only if $s_{i}s_{j}(x_{i}-x_{j})=0$ for all $1\leq i,j\leq m$.
Proof.
The result relies on the following inequality:
$t(x_{i})h(x_{j})+t(x_{j})h(x_{i})\leq t(x_{i})h(x_{i})+t(x_{j})h(x_{j}),~{}~{}\forall 1\leq i,j\leq m.$
In fact, the above is equivalent to
$(t(x_{i})-t(x_{j}))(h(x_{j})-h(x_{i}))\leq 0$
which is true by assumptions on $t(x)$ and $h(x)$. The equality holds if and only if $x_{j}=x_{i}$. Thus, we have
$\left(\sum_{j=1}^{m}s_{j}t(x_{j})\right)\left(\sum_{j=1}^{m}s_{j}h(x_{j})\right)$ $=\sum_{j=1}^{m}s_{j}^{2}t(x_{j})h(x_{j})+\sum_{1\leq i<j\leq n}s_{i}s_{j}\left(t(x_{i})h(x_{j})+t(x_{j})h(x_{i})\right)$
$\leq\sum_{j=1}^{m}s_{j}^{2}t(x_{j})h(x_{j})+\sum_{1\leq i<j\leq n}s_{i}s_{j}\left(t(x_{i})h(x_{i})+t(x_{j})h(x_{j})\right)$
$=\sum_{j=1}^{m}s_{j}(s_{1}+\dots+s_{m})t(x_{j})h(x_{j})$
$=\sum_{j=1}^{m}s_{j}t(x_{j})h(x_{j}),$
and the equality holds if and only if $s_{i}s_{j}(x_{i}-x_{j})=0$ for all $1\leq i,j\leq m$. ∎
Theorem 5.1.
At the $k$th step of Lawson’s iteration (Algorithm 1), for $\boldsymbol{w}^{(k)}\in{\cal S}$, assume $d_{2}(\boldsymbol{w}^{(k)})$ is a simple eigenvalue of the matrix pencil $(A_{\boldsymbol{w}^{(k)}},B_{\boldsymbol{w}^{(k)}})$ given in (2.13), and (A1) and (A2) hold. Then
  • i)
    there is a $\beta_{0}>0$ so that for any $\beta\in(0,\beta_{0})$, Lawson’s iteration gives $d_{2}(\boldsymbol{w}^{(k+1)})\geq d_{2}(\boldsymbol{w}^{(k)})$;
  • ii)
    for any sufficiently $\beta>0$,
    $d_{2}(\boldsymbol{w}^{(k+1)})=d_{2}(\boldsymbol{w}^{(k)})\Longrightarrow w_{j}^{(k)}r_{j}^{(k)}\left(r_{j}^{(k)}-c\right)=0,~{}\forall 1\leq j\leq m,$ (5.2)
    where $r_{j}^{(k)}=|f_{j}-\xi^{(k)}(x_{j})|$ and the constant $c$ satisfies $\sqrt{d_{2}(\boldsymbol{w}^{(k)})}\leq c\leq\|\boldsymbol{f}-\xi^{(k)}(\boldsymbol{x})\|_{\infty};$
  • iii)
    for the item ii), when ${\cal I}_{\boldsymbol{w}^{(k)}}:=\{j\in[m]|w_{j}^{(k)}\neq 0\}={\cal I}_{\boldsymbol{w}^{(k+1)}}$, then $c=\sqrt{d_{2}(\boldsymbol{w}^{(k)})}.$ In this case, if, additionally, $c\geq\max_{j\not\in{\cal I}_{\boldsymbol{w}^{(k)}}}r_{j}^{(k)}$, then $\xi^{(k)}$ is the minimax approximant of (1.2).
Proof.
For simplicity, we adopt the notation in the proof of Theorem 3.1 by omitting the subscript, and we assume that $d(\boldsymbol{w})>0$ (the conclusions are trivial when $d(\boldsymbol{w})=0$ as $w_{j}r_{j}=0,~{}\forall j\in[m]$). The idea for the proof is to express and estimate the lower bound in (3.2) using the real parameter $\beta$ around $\beta=0$.
Define $W(\beta)={\rm diag}(w_{1}(\beta),\dots,w_{m}(\beta))$ with
$w_{j}(\beta)=\frac{w_{j}r_{j}^{\beta}}{\gamma_{\beta}}=\frac{w_{j}|f_{j}-\xi(x_{j})|^{\beta}}{\sum_{j=1}^{m}w_{j}|f_{j}-\xi(x_{j})|^{\beta}}$
for which we have $w_{j}(0)=w_{j}$ and
$w_{j}^{\prime}(0)=w_{j}\log r_{j}-w_{j}\sum_{i=1}^{m}w_{i}\log r_{i},\quad\forall j\in{\cal I}_{\widetilde{\boldsymbol{w}}}.$
In (iii) of Proposition 2.1, based on assumptions (A1) and (A2), it is true that the matrices $Q_{p}(\beta),R_{p}(\beta),Q_{q}(\beta)$ and $R_{q}(\beta)$ in QR factorizations of $\sqrt{W(\beta)}\Phi=Q_{q}(\beta)R_{q}(\beta)$ and $\sqrt{W(\beta)}\Psi=Q_{p}(\beta)R_{p}(\beta)$, are all locally differentiable for sufficiently small $\beta$. Moreover, as $d_{2}(\boldsymbol{w}(0))$ is a simple eigenvalue of the matrix pencil $(A_{\boldsymbol{w}(0)},B_{\boldsymbol{w}(0)})$ given in (2.13), by (2.14), it is also a simple eigenvalue of the Hermitian matrix $S_{F}(\beta)-S_{qp}(\beta)S_{qp}(\beta)^{\operatorname{H}}$ at $\beta=0$. Thus, the continuity of eigenvalues implies that the smallest eigenvalue $d_{2}(\boldsymbol{w}(\beta))$ is a simple eigenvalue of the Hermitian $S_{F}(\beta)-S_{qp}(\beta)S_{qp}(\beta)^{\operatorname{H}}$ for any sufficiently small $\beta\in\mathbb{R}$. Alternatively, we can say that the eigenspace spanned by the unit-norm eigenvector $R_{q}(\beta)\boldsymbol{b}(\boldsymbol{w}(\beta))$ of $S_{F}(\beta)-S_{qp}(\beta)S_{qp}(\beta)^{\operatorname{H}}$ corresponding to $d_{2}(\boldsymbol{w}(\beta))$ is one-dimensional. Based on [20, Chapter 2.6.2], there is a continuously differentiable normalized eigenvector $R_{q}(\beta)\boldsymbol{b}(\boldsymbol{w}(\beta))$ with respect to $\beta$ around $\beta=0$. Moreover, noting that $\boldsymbol{q}(\beta)=\Phi\boldsymbol{b}(\beta)=\Phi(R_{q}(\beta))^{-1}R_{q}(\beta)\boldsymbol{b}(\beta)$, and also that $|\boldsymbol{q}(\beta)^{\operatorname{H}}W(0)\boldsymbol{q}(0)|$ does not change for different choices of a unit-norm eigenvector $R_{q}(\beta)\boldsymbol{b}(\boldsymbol{w}(\beta))$, in the following discussion, we can assume $\boldsymbol{q}(\beta)$ is continuously differentiable with respect to $\beta$ around $\beta=0$. Hence,
$\boldsymbol{q}(\beta)=\boldsymbol{q}(0)+\beta\boldsymbol{q}^{\prime}(0)+O(\beta^{2}).$
Since
$1=\boldsymbol{q}(\beta)^{\operatorname{H}}W(\beta)\boldsymbol{q}(\beta)\Longrightarrow{\rm Re}(\boldsymbol{q}^{\prime}(0)^{\operatorname{H}}W(0)\boldsymbol{q}(0))=-\frac{1}{2}\boldsymbol{q}(0)^{\operatorname{H}}W^{\prime}(0)\boldsymbol{q}(0),$
for any sufficiently small $\beta\in\mathbb{R}$, we have
$\left|\boldsymbol{q}(\beta)^{\operatorname{H}}W(0)\boldsymbol{q}(0)\right|$
$\geq\left|{\rm Re}(\boldsymbol{q}(\beta)^{\operatorname{H}}W(0)\boldsymbol{q}(0))\right|$
$=\left|\boldsymbol{q}(0)^{\operatorname{H}}W(0)\boldsymbol{q}(0)+\beta{\rm Re}(\boldsymbol{q}^{\prime}(0)^{\operatorname{H}}W(0)\boldsymbol{q}(0))\right|+O(\beta^{2})$
$=\left|1+\beta{\rm Re}(\boldsymbol{q}^{\prime}(0)^{\operatorname{H}}W(0)\boldsymbol{q}(0))\right|+O(\beta^{2})$
$=\left|1-\frac{\beta}{2}\boldsymbol{q}(0)^{\operatorname{H}}W^{\prime}(0)\boldsymbol{q}(0)\right|+O(\beta^{2})$
$=1-\frac{\beta}{2}\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}|q_{j}|^{2}w_{j}\log r_{j}-\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}|q_{j}|^{2}w_{j}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{i}\log r_{i}\right)\right)+O(\beta^{2})$
$=1-\frac{\beta}{2}\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}|q_{j}|^{2}w_{j}\log r_{j}-\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{i}\log r_{i}\right)+O(\beta^{2}).$
With this, we can write the lower bound in (3.2) as
^$\widehat{\ell}(\beta):$ $=\frac{|\boldsymbol{q}(\beta)^{\operatorname{H}}W(0)\boldsymbol{q}(0)|}{\gamma_{\beta}~{}\zeta_{\beta}}$
$\geq\underbrace{\frac{1-\frac{\beta}{2}\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}|q_{j}|^{2}w_{j}\log r_{j}-\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{i}\log r_{i}\right)}{\sqrt{\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}\right)}}}_{=:\ell(\beta)}+O(\beta^{2})$
$=:\ell(\beta)+O(\beta^{2})$
locally at $\beta=0$.
For $\ell(\beta)$, by calculation, we have (with $q_{j}=q_{j}(0),~{}r_{j}=r_{j}(0)$ and $w_{j}=w_{j}(0)$)
$\ell^{\prime}(0)$ $=\frac{1}{2\sqrt{(d_{2}(\boldsymbol{w}(0)))^{3}}}\left[\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2}\log r_{j}-\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}\log r_{j}\right)\right]$
$\geq 0$
where the last inequality is obtained by applying Lemma 5.1 with $s_{j}=w_{j}|q_{j}|^{2}$, $x_{j}=r_{j}$, $t(x)=x^{2}$ and $h(x)=\log x$ on the interval $(0,\infty)$. Furthermore, if there is a pair $(i,j)$ so that ${i\in{\cal I}_{\widetilde{\boldsymbol{w}}}}$, ${j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}$ and $w_{i}w_{j}(r_{i}-r_{j})\neq 0$, then we have $\ell^{\prime}(0)>0$ by Lemma 5.1. In that case, we know that there is a $\beta_{0}>0$ such that $\ell^{\prime}(\beta)>\frac{1}{2}\ell^{\prime}(0)$ and $\frac{\beta}{2}\ell^{\prime}(0)+O(\beta^{2})>0$ for any $\beta\in(0,\beta_{0})$, implying
[…]
This shows that, if there is a pair $(i,j)$ so that ${i,j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}$ and $w_{i}w_{j}(r_{i}-r_{j})\neq 0$, then a sufficiently small $\beta>0$ leads to $d_{2}(\boldsymbol{w}(\beta))>d_{2}(\boldsymbol{w}(0))$; in other words, for a sufficiently small $\beta>0$,
$d_{2}(\boldsymbol{w}(\beta))=d_{2}(\boldsymbol{w}(0))\Longrightarrow w_{j}w_{i}(r_{i}-r_{j})=0,~{}\forall{i,j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}\Longrightarrow r_{j}={\rm a~{}constant~{}}c,~{}\forall{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}.$ (5.3)
Observe that $c\leq\max_{j\in[m]}r_{j}=\|\boldsymbol{f}-\xi(\boldsymbol{x})\|_{\infty}$ and also
$d_{2}(\boldsymbol{w}(0))=\sum_{j=1}^{m}w_{j}|q_{j}r_{j}|^{2}=\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|r_{j}|^{2}|q_{j}|^{2}=c^{2}\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}\leq c^{2}\sum_{j=1}^{m}w_{j}|q_{j}|^{2}=c^{2},$ (5.4)
giving $\sqrt{d_{2}(\boldsymbol{w}(0))}\leq c\leq\|\boldsymbol{f}-\xi(\boldsymbol{x})\|_{\infty}.$ This proves items i) and ii).
For iii), we note that ${\cal I}_{\boldsymbol{w}}={\cal I}_{\widetilde{\boldsymbol{w}}}$ implies $r_{j}\neq 0$ and $r_{j}=c,~{}\forall j\in{\cal I}_{\boldsymbol{w}}$ (see (5.3)). Thus from (5.4),
$d_{2}(\boldsymbol{w}(0))=\sum_{j=1}^{m}w_{j}|q_{j}r_{j}|^{2}=\sum_{j\in{\cal I}_{\boldsymbol{w}}}w_{j}|r_{j}|^{2}|q_{j}|^{2}=c^{2}\sum_{j\in{\cal I}_{\boldsymbol{w}}}w_{j}|q_{j}|^{2}=c^{2}\sum_{j=1}^{m}w_{j}|q_{j}|^{2}=c^{2}.$
If we additionally have $c\geq\max_{j\not\in{\cal I}_{\boldsymbol{w}}}r_{j}$, then $c=\sqrt{d_{2}(\boldsymbol{w}(0))}=\|\boldsymbol{f}-\xi(\boldsymbol{x})\|_{\infty}$, which, according to Theorem 2.3 ([43, Theorem 4.3]), implies that Ruttan’s sufficient condition (or, equivalently, the strong duality) for the minimax approximation is satisfied, and $\xi^{(k)}$ is the minimax approximant of (1.2). In this case, the result in (5.2) can be written as
$w_{j}\left(r_{j}-\|\boldsymbol{f}-\xi(\boldsymbol{x})\|_{\infty}\right)=0,~{}\forall j\in[m],$
which is the complementary slackness property in (2.12). ∎
Remark 5.1.
We remark first that the conclusion $w_{j}^{(k)}r_{j}^{(k)}\left(r_{j}^{(k)}-c\right)=0,~{}\forall j\in[m]$ in (5.2) is a certain complementary slackness. Indeed, it says that for node $x_{j}$ with $w_{j}^{(k)}>0$, either $r_{j}^{(k)}=0$ or $r_{j}^{(k)}=c$; that is, any node associated with positive weight either is an interpolation point or has the error $c$.
According to our proof for Theorem 5.1, the Lawson exponent $\beta_{0}>0$ in Theorem 5.1 is dependent on how far the continuously differentiable normalized eigenvector $R_{q}(\beta)\boldsymbol{b}(\boldsymbol{w}(\beta))$ of the Hermitian matrix $S_{F}(\beta)-S_{qp}(\beta)S_{qp}(\beta)^{\operatorname{H}}$ can be extended from $\beta=0$. Explicit formulation for $\beta_{0}$ is hard, but intuitively, the larger gap between the smallest eigenvalue $d_{2}(\boldsymbol{w})$ and the next eigenvalue is, the larger $\beta_{0}$ is. Currently, the proof does not guarantee $\beta_{0}>1$, which, however, performs always well and is a recommended value in practice [43].
6 Conclusions
In this paper, we have established theoretical guarantees for Lawson’s iteration in solving both the linear and rational minimax approximations. For the rational minimax approximation, our results indicate that, generically, a small Lawson exponent $\beta>0$ leads to the monotonic convergence, and also reveal some limit properties of Lawson’s iteration. These theoretical guarantees, on the one hand, explain some numerical behaviors (for example, “taking $\beta$ to be smaller makes the algorithm much more robust[11]), and on the other hand, provide more insights on why and how Lawson’s updating scheme works. These theoretical results lay a solid ground for this version of Lawson’s iteration.