### v3

# A convergence analysis of Lawson’s iteration for computing polynomial and rational minimax approximations

Abstract
Lawson’s iteration is a classical and effective method for solving the linear (polynomial) minimax approximation problem in the complex plane. Extension of Lawson’s iteration for the rational minimax approximation problem with both computationally high efficiency and theoretical guarantee is challenging. A recent work [L.-H. Zhang, L. Yang, W. H. Yang and Y.-N. Zhang, A convex dual problem for the rational minimax approximation and Lawson’s iteration, Math. Comp., 94(2025), 2457–2494.] reveals that Lawson’s iteration can be viewed as a method for solving the dual problem of the original rational minimax approximation problem, and a new type of Lawson’s iteration, namely, d-Lawson, was proposed, which reduces to the classical Lawson’s iteration for the linear minimax approximation problem. For the rational case, such a dual problem is guaranteed to obtain the original minimax solution under Ruttan’s sufficient condition, and numerically, d-Lawson was observed to converge monotonically with respect to the dual objective function. In this paper, we present a theoretical convergence analysis of d-Lawson for both the linear and rational minimax approximation problems. In particular, we show that
  • (i)
    for the linear minimax approximation problem, $\beta=1$ is a near-optimal Lawson exponent in Lawson’s iteration, and
  • (ii)
    for the rational minimax approximation problem, under certain conditions, d-Lawson converges monotonically with respect to the dual objective function for any sufficiently small $\beta>0$, and the limiting approximant satisfies the complementary slackness condition: any node associated with positive weight either is an interpolation point or has a constant error.
Key words. Rational minimax approximation, Lawson algorithm, Ruttan’s optimality condition, Dual programming, Convergence analysis
AMS subject classifications. 41A50, 41A20, 65D15, 90C46
1 Introduction
Computing the polynomial and/or rational minimax (also known as Chebyshev or best) approximations of a continuous complex-valued function $f$ over a given compact set $\Omega$ in the complex plane $\mathbb{C}$ is a classical problem in approximation theory [42]. In practical applications, rational approximations are useful in various areas, including function approximations [11, 44], computing conformal mappings [18, 22, 43], solving partial differential equations [5, 8, 17, 19, 32, 41, 45, 51], model order reduction [2, 9, 20], and signal processing [1, 10, 15, 16, 23, 40, 46, 47]; see [31] for some recent applications. In many of these applications, only discretized data samples are available; even for a continuum domain $\Omega$ enclosed by a simple Jordan curve in which $f$ is analytic, by the maximum modulus principle, we can first sample $f$ on the boundary of $\Omega$, and then compute the rational/polynomial minimax approximant of $f$ through solving a discrete rational/polynomial minimax problem. For these cases, denote by $\{(x_{j},f_{j})\}_{j=1}^{m}$ the sampled data from $f_{j}=f(x_{j})\in\mathbb{C}$ ($x_{j}\in\Omega$) over distinct nodes ${\cal X}=\{x_{j}\}_{j=1}^{m}$, and denote by $\mathbb{P}_{n}$ the set of complex polynomials with degree less than or equal to $n$. We consider the following discrete rational approximation problem
$\inf_{\xi=p/q\in\mathscr{R}_{(n_{1},n_{2})}}\|\boldsymbol{f}-\xi(\boldsymbol{x})\|_{\infty},$ (1.1)
where $\mathscr{R}_{(n_{1},n_{2})}:=\{{p}/{q}|p\in\mathbb{P}_{n_{1}},~0\not\equiv q\in\mathbb{P}_{n_{2}}\}$, $\boldsymbol{f}=[f_{1},\dots,f_{m}]^{\operatorname{T}}\in\mathbb{C}^{m}~(n_{1}+n_{2}+2\leq m)$, $\boldsymbol{x}=[x_{1},\dots,x_{m}]^{\operatorname{T}}\in\mathbb{C}^{m}$, $\xi(\boldsymbol{x})=[\xi(x_{1}),\dots,\xi(x_{m})]^{\operatorname{T}}\in\mathbb{C}^{m}$, and
$\|\boldsymbol{f}-\xi(\boldsymbol{x})\|_{\infty}:=\max_{1\leq j\leq m}\left|f_{j}-\frac{p(x_{j})}{q(x_{j})}\right|.$
In case when the infimum of (1.1) is attainable, we call the function $\xi^{*}={p^{*}}/{q^{*}}\in\mathscr{R}_{(n_{1},n_{2})}$ from
${p^{*}}/{q^{*}}\in\arg\min_{\xi=p/q\in\mathscr{R}_{(n_{1},n_{2})}}\|\boldsymbol{f}-\xi(\boldsymbol{x})\|_{\infty},$ (1.2)
the rational minimax approximant [42] of $f(x)$ over $\cal X$}.
In general, computing the discrete rational minimax approximation is much more challenging than the polynomial (i.e., $n_{2}=0$) minimax problem. Indeed, for the polynomial case, it is guaranteed that there is a unique minimax approximant which can be characterized by a necessary and sufficient condition (e.g., [35] and [48, Theorem 2.1]); for the rational case, the infimum of (1.1) may not be achievable, and even if it is, there may be multiple minimax approximants [30, 42]. Furthermore, local best approximants may exist [30, 42]. Necessary conditions for the rational minimax approximant have been developed in, e.g., [13, 21, 36, 37, 39, 48, 49, 50], and Ruttan [36, Theorem 2.1] contributes a sufficient condition.
Lawson’s iteration [27] is a classical and effective method for computing the discrete linear (polynomial) minimax approximant. The idea of Lawson is to approximate the minimax polynomial $p^{*}$ (i.e., $n_{2}=0$ in (1.2)) by a sequence of polynomials $\{p^{(k)}\}$, each as the solution of the weighted least-squares problem:
$p^{(k)}=\operatorname*{argmin}_{p\in\mathbb{P}_{n}}\sum_{j=1}^{m}w_{j}^{(k)}|f_{j}-p(x_{j})|^{2},$
where $\boldsymbol{w}^{(k)}=[w_{1}^{(k)},\dots,w_{m}^{(k)}]^{\operatorname{T}}\in{\cal S}$ is the weight vector at the $k$th iteration in the probability simplex:
${\cal S}:=\{\boldsymbol{w}=[w_{1},\dots,w_{m}]^{\operatorname{T}}\in\mathbb{R}^{m}:\boldsymbol{w}\geq 0~{\rm and}~\boldsymbol{w}^{\operatorname{T}}\boldsymbol{e}=1\},~~\boldsymbol{e}=[1,1,\dots,1]^{\operatorname{T}}.$
At the $k$th step in Lawson’s iteration, it updates the element of the weight vector as
$w_{j}^{(k+1)}=\frac{w_{j}^{(k)}\left|f_{j}-p^{(k)}(x_{j})\right|^{\beta}}{\sum_{i}w_{i}^{(k)}\left|f_{i}-p^{(k)}(x_{i})\right|^{\beta}},~~\forall j,$ (1.3)
where $\beta>0$ is the so-called Lawson exponent and is originally set as $\beta=1$ (see e.g., [6, 27]). Due to its relation with weighted least-squares problems, Lawson’s iteration is an iteratively reweighted least-squares (IRLS) iteration. Convergence analysis and some variants have been discussed (e.g., [3, 6, 7, 24, 28, 38, 52]).
However, extension of Lawson’s iteration for the rational minimax approximation problem with both computationally high efficiency and theoretical guarantee is nontrivial and challenging. Particularly, for computing the minimax rational solution $\xi^{*}=p^{*}/q^{*}$ of (1.2), two difficulties related with a basic Lawson’s step (1.3) are: how to define a suitable approximation $\xi^{(k)}=p^{(k)}/q^{(k)}$ associated with the current weight vector $\boldsymbol{w}^{(k)}$, and how to choose a suitable Lawson exponent $\beta$ for convergence?
Some versions of Lawson’s iteration have been discussed for the rational minimax approximation problem (1.1). For example, the Loeb algorithm [28] (the same method was also proposed in [38] known as the SK iteration by Sanathanan and Koerner) uses the reciprocal of the current denominator $q^{(k)}(x_{j})$ as weights and compute the approximation $\xi^{(k)}=p^{(k)}/q^{(k)}$ from a weighted linearization associated with the current weights. A recent work [24] further improves the basic SK iteration by proposing a stabilized SK iteration. Another remarkable work in the rational approximation literature is the adaptive Antoulas-Anderson (AAA) algorithm [29] and its extension, the AAA-Lawson algorithm [14, 30, 31]. AAA represents the rational approximation in barycentric form and selects the associated support points iteratively in an adaptive way for stability; in AAA-Lawson [14, 30, 31], the algorithm further introduces weights updated according to [14, Equ. (8.5)]
$w_{j}^{(k+1)}=\frac{w_{j}^{(k)}\left|f_{j}-\xi_{\tiny\rm AAA-Lawson}^{(k)}(x_{j})\right|^{\beta}}{\sum_{i}w_{i}^{(k)}\left|f_{i}-\xi_{\tiny\rm AAA-Lawson}^{(k)}(x_{i})\right|^{\beta}},~~\forall j,$ (1.4)
where $\xi_{\tiny\rm AAA-Lawson}^{(k)}$ is from a weighted linearization [14, Equ. (8.4)] associated with the current weight vector $\boldsymbol{w}^{(k)}$ (see also (2.19)). Other versions of Lawson’s iteration can be found in [3, 7]. However, to our best knowledge, no convergence guarantee has been established for these versions of Lawson’s iteration in theory, and in some cases, the computed rational approximants can be local best or near-best [14, 30]. Indeed, as remarked for AAA-Lawson in [14] that “its convergence is far from understood, and even when it does converge, the rate is slow (linear at best)”, and “convergence analysis appears to be highly nontrivial”.
In this paper, we shall establish the convergence of a version of Lawson’s iteration (Algorithm 1), namely, d-Lawson, proposed recently in [54]. In the same Lawson’s updating fashion for the weights $w_{j}^{(k+1)}$ given in (1.4), the main difference between d-Lawson and AAA-Lawson lies in defining the associated approximant $\xi^{(k)}=p^{(k)}/q^{(k)}$, and we shall clarify this difference in detail in section 2.5. For d-Lawson, [54] reveals that it can be viewed as a method for solving the dual problem $\max_{\boldsymbol{w}\in{\cal S}}d_{2}(\boldsymbol{w})$ of the original rational minimax approximation problem. The weight $w_{j}$ is the corresponding dual variable associated with the node $x_{j}$. Such a dual problem is guaranteed to obtain the original minimax solution $\xi^{*}$ under Ruttan’s sufficient condition ([36, Theorem 2.1]; see also [25, Theorem 2] and [39, Theorem 3]); moreover, numerically, d-Lawson was observed to converge monotonically [54] with respect to the dual objective function $d_{2}(\boldsymbol{w})$. The framework for handling the rational minimax approximation problem (1.1) in [54] can be well described by Figure 1.1 (see [54, Figure 1]). It should be pointed out that d-Lawson reduces to the classical Lawson’s iteration [27] for the linear minimax approximation problem if $n_{2}=0$.
Primal problem: (Chebyshevappr.){(Chebyshev~appr.)~} $\rm min-max$} Dual problem $\rm(max-min)$}: $\max_{\boldsymbol{w}\in{\cal S}}d_{2}(\boldsymbol{w})$ $\rm Lawson^{\prime}s~iteration$} solve the dualweak $\checkmarkstrong duality$ underRuttan’s conditionLagrange duality on a linearization
Figure 1.1: Framework [54, Figure 1] of the dual programming and Lawson’s iteration for the rational minimax approximation of (1.2).
For this version d-Lawson of Lawson’s iteration (Algorithm 1), we shall contribute theoretical convergence analysis for both the linear and rational minimax approximation problems. In particular, we show that
  • (i)
    for the linear minimax approximation problem, $\beta=1$ is a near-optimal Lawson exponent in Lawson’s iteration (1.3), and
  • (ii)
    for the rational minimax approximation problem, under certain conditions, d-Lawson converges monotonically with respect to the dual objective function $d_{2}(\boldsymbol{w})$ for any sufficiently small $\beta>0$ in (1.4); moreover, the convergent pair $(\boldsymbol{w},\xi)$ satisfies the following complementary slackness conditions:
    $w_{j}|f_{j}-\xi(x_{j})|\left(|f_{j}-\xi(x_{j})|-c\right)=0,\quad\forall j=1,2,\dots,m,$ (1.5)
    for some constant $c$. The complementary slackness means that any node associated with positive weight either is an interpolation point or has a constant error.
These theoretical findings lay a solid ground for d-Lawson (Algorithm 1).
We organize the paper as follows. In section 2, we first introduce the dual problem associated with a linearization of the rational minimax approximation problem (1.1) using Lagrange duality [54]. These results are mainly from [54]. We shall introduce the concept of strong duality and explore its connection with Ruttan’s sufficient condition; an optimality condition for computing the dual objective function and d-Lawson [54] (Algorithm 1) for solving the dual problem will be presented; furthermore, a roadmap of our convergence analysis and a comparison between AAA-Lawson and d-Lawson are given in this section. In section 3, we shall establish an important lower bound for the dual objective function value $d_{2}(\boldsymbol{w}^{(k+1)})$ based on the information at the $k$th iteration. Relying on this lower bound, we prove in section 4 that for the linear minimax approximation problem, there is a $\beta_{0}>1$ so that for any $\beta\in(0,\beta_{0})$, Lawson’s iteration converges monotonically, and $\beta=1$ is the near-optimal Lawson exponent. In section 5, we will use the lower bound to show that monotonic convergence of d-Lawson occurs generically for any sufficiently small $\beta>0$, and furthermore, we shall prove the complementary slackness whenever $d_{2}(\boldsymbol{w}^{(k+1)})=d_{2}(\boldsymbol{w}^{(k)})$. Finally concluding remarks are drawn in section 6.
Notation. We follow the notation in [54] in this paper. The imaginary unit is ${\tt i}=\sqrt{-1}$, and for $\mu=\mu^{\tt r}+{\tt i}\mu^{\tt i}\in\mathbb{C}$, we denote its modulus $|\mu|=\sqrt{(\mu^{\tt r})^{2}+(\mu^{\tt i})^{2}}$ and its conjugate $\bar{\mu}=\mu^{\tt r}-{\tt i}\mu^{\tt i}$, where ${\rm Re}(\mu)=\mu^{\tt r}\in\mathbb{R}$ and ${\rm Im}(\mu)=\mu^{\tt i}\in\mathbb{R}$ are the real and imaginary part of $\mu$, respectively. Bold lower case letters are used to represent column vectors, and ${\mathbb{C}}^{n\times m}$ (resp. ${\mathbb{R}}^{n\times m}$) stands for the set of all $n\times m$ complex (resp. real) matrices, with the identity $I_{n}\equiv[\boldsymbol{e}_{1},\boldsymbol{e}_{2},\dots,\boldsymbol{e}_{n}]\in\mathbb{R}^{n\times n}$, where $\boldsymbol{e}_{i}$ is its $i$th column with $i\in\{1,2,\dots,n\}$. For a vector $\boldsymbol{x}\in\mathbb{C}^{n}$, ${\rm diag}(\boldsymbol{x})={\rm diag}(x_{1},\dots,x_{n})$ is the diagonal matrix, and $\|\boldsymbol{x}\|_{\alpha}=(\sum_{j=1}^{n}|x_{j}|^{\alpha})^{\frac{1}{\alpha}}$ is the vector $\alpha$-norm ($\alpha\geq 1$) of $\boldsymbol{x}$. For $\boldsymbol{x},\boldsymbol{y}\in\mathbb{C}^{n}$ with $y_{j}\neq 0,~1\leq j\leq n$, we define $\boldsymbol{x}./\boldsymbol{y}=[x_{1}/y_{1},\dots,x_{n}/y_{n}]^{\operatorname{T}}.$ For a matrix $A\in\mathbb{C}^{m\times n}$, ${\rm span}(A)$ represents the column space of $A$; $A^{\operatorname{H}}$ (resp. $A^{\operatorname{T}}$) and $A^{{\dagger}}$ are the conjugate transpose (resp. transpose) and the Moore-Penrose inverse of $A$, respectively. We also adopt MATLAB-like convention to represent the sub-matrix $A({\cal I}_{1},{\cal I}_{2})$ of $A$, consisting of intersections of rows and columns indexed by ${\cal I}_{1}\subseteq\{1,2,\dots,m\}$ and ${\cal I}_{2}\subseteq\{1,2,\dots,n\}$, respectively.
2 A dual problem and Lawson’s iteration
First, let
$\mathbb{P}_{n_{1}}={\rm span}(\psi_{0}(x),\dots,\psi_{n_{1}}(x))~{\rm and}~\mathbb{P}_{n_{2}}={\rm span}(\phi_{0}(x),\dots,\phi_{n_{2}}(x))$
be the chosen bases for the numerator and denominator polynomial spaces, respectively, and we write $p/q\in\mathscr{R}_{(n_{1},n_{2})}$ as
$\frac{p(x)}{q(x)}=\frac{[\psi_{0}(x),\dots,\psi_{n_{1}}(x)]\boldsymbol{a}}{[\phi_{0}(x),\dots,\phi_{n_{2}}(x)]\boldsymbol{b}},~~\mbox{for ~some~}\boldsymbol{a}\in\mathbb{C}^{n_{1}+1},~\boldsymbol{b}\in\mathbb{C}^{n_{2}+1}.$ (2.1)
For the given ${\cal X}=\{x_{j}\}_{j=1}^{m}$ with $|{\cal X}|=m$, we have the basis matrix for $p\in\mathbb{P}_{n_{1}}$:
$\Psi=\Psi(x_{1},\dots,x_{m};n_{1}):=\left[\begin{array}[]{cccc}\psi_{0}(x_{1})&\psi_{1}(x_{1})&\cdots&\psi_{n_{1}}(x_{1})\\ \psi_{0}(x_{2})&\psi_{1}(x_{2})&\cdots&\psi_{n_{1}}(x_{2})\\ \vdots&\cdots&\cdots&\vdots\\ \psi_{0}(x_{m})&\psi_{1}(x_{m})&\cdots&\psi_{n_{1}}(x_{m})\end{array}\right],~\Psi_{i,j}=\psi_{j-1}(x_{i}),$
and analogously, we have $\Phi=\Phi(x_{1},\dots,x_{m};n_{2})=[\phi_{j-1}(x_{i})]\in\mathbb{C}^{m\times(n_{2}+1)}$.
For a given irreducible $\xi(x)=p(x)/q(x)\in\mathscr{R}_{(n_{1},n_{2})}$, if $|\xi(x)|$ is bounded for any $x\in{\cal X}$, then it is easy to see that $q(x)\neq 0$. We define the maximum error
$e(\xi):=\max_{x\in{\cal X}}|f(x)-\xi(x)|=\|\boldsymbol{f}-\xi(\boldsymbol{x})\|_{\infty}.$ (2.2)
The defect of an irreducible $\xi(x)=p(x)/q(x)\in\mathscr{R}_{(n_{1},n_{2})}$ is
$\upsilon(p,q):=\min(n_{1}-\deg(p),n_{2}-\deg(q)),$ (2.3)
where $\deg(p)$ and $\deg(q)$ are the degrees of $p$ and $q$, respectively. When $\upsilon(p,q)=0$, we say $\xi(x)=p(x)/q(x)$ is non-degenerate. As a necessary condition for the irreducible minimax rational approximant $\xi^{*}=p^{*}/q^{*}$ of (1.2), the following result states that the number of reference points (aka the extreme points), i.e., nodes $x_{j}\in{\cal X}$ that achieve
$\left|f_{j}-\frac{p^{*}(x_{j})}{q^{*}(x_{j})}\right|=e(\xi^{*}),$
is at least $n_{1}+n_{2}+2-\upsilon(p^{*},q^{*})$.
Theorem 2.1.
([21, Theorem 2.5]) Given $m\geq{n_{1}}+n_{2}+2$ distinct nodes ${\cal X}=\{x_{j}\}_{j=1}^{m}$ on $\Omega$, suppose $\xi^{*}=p^{*}/q^{*}\in\mathscr{R}_{(n_{1},n_{2})}$ is an irreducible rational polynomial and denote the extremal set ${\cal X}_{e}(\xi^{*})\subseteq{\cal X}$ by
${\cal X}_{e}(\xi^{*}):=\left\{x_{j}\in{\cal X}:\left|f_{j}-\frac{p^{*}(x_{j})}{q^{*}(x_{j})}\right|=e(\xi^{*})\right\}.$ (2.4)
If $\xi^{*}$ is a solution to (1.2) with $\eta_{\infty}=\|\boldsymbol{f}-\xi^{*}(\boldsymbol{x})\|_{\infty},$ then the cardinality $|{\cal X}_{e}(\xi^{*})|\geq n_{1}+n_{2}+2-\upsilon(p^{*},q^{*})$; that is, ${\cal X}_{e}(\xi^{*})$ contains at least $n_{1}+n_{2}+2-\upsilon(p^{*},q^{*})$ nodes.
2.1 A linearization
In [54], by introducing a real variable $\eta$, the original minimax problem (1.1) is transformed into the following linearization:
$\inf_{\eta\in\mathbb{R},~p\in\mathbb{P}_{n_{1}},~{0\not\equiv q}\in\mathbb{P}_{n_{2}}}\eta$
$s.t.,$~ $|f_{j}q(x_{j})-p(x_{j})|^{{{2}}}\leq\eta|q(x_{j})|^{{{2}}},~~\forall j=1,2,\dots,m.$ (2.5)
Unlike the original bi-level min-max problem (1.1), problem (2.5) is a single-level optimization, whose infimum is also attainable.
Theorem 2.2.
Given $m\geq{n_{1}}+n_{2}+2$ distinct nodes ${\cal X}=\{x_{j}\}_{j=1}^{m}$ on $\Omega\subset\mathbb{C}$, let $\eta_{2}$ be the infimum of (2.5). Then $\eta_{2}$ is attainable by a pair $(p,q)$ with $p\in\mathbb{P}_{n_{1}}$ and $0\not\equiv q\in\mathbb{P}_{n_{2}}$.
Proof.
For any $0\not\equiv q\in\mathbb{P}_{n_{2}}$, as $m\geq{n_{1}}+n_{2}+2$, there exists $x_{j}\in\cal{X}$ with $q(x_{j})\neq 0$. This implies that $0\leq\sqrt{\eta_{2}}\leq f_{\max}:=\max_{1\leq j\leq m}|f_{j}|$ where $\sqrt{\eta_{2}}\leq f_{\max}$ is obtained by choosing a feasible solution ${\eta}=(f_{\max})^{2},~p\equiv 0$ and $q\equiv 1$. Suppose $\{(\eta^{(k)},p^{(k)},q^{(k)})\}$ is the feasible sequence so that $p^{(k)}\in\mathbb{P}_{n_{1}}$, $0\not\equiv q^{(k)}\in\mathbb{P}_{n_{2}}$ and $\eta^{(k)}\rightarrow\eta_{2}$ as $k\rightarrow\infty$. Noting that $\{(\eta^{(k)},\tau^{(k)}p^{(k)},\tau^{(k)}q^{(k)})\}$ is also feasible for (2.5) for any $\tau^{(k)}\neq 0$, we then can choose $\tau^{(k)}$ so that the coefficient vector $\boldsymbol{b}^{(k)}=[b_{0}^{(k)},\dots,b_{n_{2}}^{(k)}]^{\operatorname{T}}\in\mathbb{C}^{n_{2}+1}$ of $\tau^{(k)}q^{(k)}(x)=\sum_{i=0}^{n_{2}}b_{i}^{(k)}x^{i}$ satisfies $\|\boldsymbol{b}^{(k)}\|_{2}=1~(\forall k\geq 0)$. Let $\tilde{p}^{(k)}:=\tau^{(k)}p^{(k)}$ and $\tilde{q}^{(k)}:=\tau^{(k)}q^{(k)}$. Thus $\{\tilde{q}^{(k)}\}$ has a convergent subsequence resulting from a convergent subsequence of $\{\boldsymbol{b}^{(k)}\}$. For simplicity of presentation, we assume that $\{\tilde{q}^{(k)}\}$ itself converges to $0\not\equiv q\in\mathbb{P}_{n_{2}}$, and $\eta^{(k)}\leq\rho_{0}~(\forall k\geq 0)$ for some $\rho_{0}>0$.
Note that $|f_{j}\tilde{q}^{(k)}(x_{j})-\tilde{p}^{(k)}(x_{j})|\leq\sqrt{\eta^{(k)}}|\tilde{q}^{(k)}(x_{j})|$ implies
$|\tilde{p}^{(k)}(x_{j})|\leq\sqrt{\eta^{(k)}}|\tilde{q}^{(k)}(x_{j})|+|f_{j}\tilde{q}^{(k)}(x_{j})|\leq(\sqrt{\rho_{0}}+{f_{\max}})\rho_{1}=:\tilde{\rho}_{1},~\forall x_{j}\in{\cal X},$
where $\rho_{1}=\sum_{i=0}^{n_{2}}|x_{J}|^{i}$ with $|x_{J}|=\max_{1\leq j\leq m}|x_{j}|$. Denoting $\tilde{p}^{(k)}(x)=\sum_{i=0}^{n_{1}}a_{i}^{(k)}x^{i}$ in the monomial basis, we can choose the first $n_{1}+1$ nodes $\{x_{j}\}_{j=1}^{n_{1}+1}$ to have
$\tilde{p}^{(k)}(x_{1}),\dots,\tilde{p}^{(k)}(x_{{n_{1}}+1})]^{\operatorname{T}}=V(x_{1},\dots,x_{n_{1}+1})[a_{0}^{(k)},\dots,a_{n_{1}}^{(k)}]^{\operatorname{T}}$
where $V(x_{1},\dots,x_{n_{1}+1})$ is the Vandermonde matrix associated with $\{x_{j}\}_{j=1}^{n_{1}+1}$. This gives
$\|[a_{0}^{(k)},\dots,a_{n_{1}}^{(k)}]\|_{2}\leq\tilde{\rho}_{1}\sqrt{n_{1}+1}\|[V(x_{1},\dots,x_{n_{1}+1})]^{-1}\|_{2},~\forall k\geq 0.$
Therefore, $\{\tilde{p}^{(k)}\}$ also has a convergent subsequence with a limit polynomial $p\in\mathbb{P}_{n_{1}}$, and consequently, $(\eta_{2},p,q)$ solves (2.5). ∎
As a further step, we shall show in Theorem 2.3 that the two infimums of (1.1) and (2.5) are consistent. Thus if (1.1) admits a solution, it can be recovered from the linearization (2.5). It is a generalization of [54, Theorem 2.1]. The next lemma is used for proving Theorem 2.3.
Lemma 2.1.
Let $1\leq s\leq\max(n_{1},n_{2})$. Given any distinct nodes $x_{1},\dots,x_{s}$ in $\mathbb{C}$ and the corresponding $f_{j}\in\mathbb{C},~j=1,\dots,s$, for any $\epsilon>0$, there is a $\frac{p}{q}\in\mathscr{R}_{(n_{1},n_{2})}$ such that $\left|f_{j}-\frac{p(x_{j})}{q(x_{j})}\right|\leq\epsilon$ and $q(x_{j})\neq 0$ $\forall j=1,\dots,s$.
Proof.
For $s\leq n_{1}$, choose $q\equiv 1$ and $p\in\mathbb{P}_{n_{1}}$ satisfying the interpolation conditions $p(x_{j})=f_{j}~(1\leq j\leq s)$; thus $\left|f_{j}-\frac{p(x_{j})}{q(x_{j})}\right|=0,~1\leq j\leq s$. For $n_{1}<s\leq n_{2}$, let $p\equiv 1$ and $q\in\mathbb{P}_{n_{2}}$ satisfying the interpolation conditions: $q(x_{j})=\frac{1}{f_{j}}$ if $f_{j}\neq 0$, and $q(x_{j})=\frac{2}{\epsilon}$ otherwise; thus $\left|f_{j}-\frac{p(x_{j})}{q(x_{j})}\right|\leq\frac{\epsilon}{2}<\epsilon,~1\leq j\leq s$. This completes the proof. ∎
Theorem 2.3.
Given $m\geq{n_{1}}+n_{2}+2$ distinct nodes ${\cal X}=\{x_{j}\}_{j=1}^{m}$ on $\Omega\subset\mathbb{C}$, let $\eta_{2}$} be the infimum of (2.5). Then $\eta_{2}=(\eta_{\infty})^{2}$. Furthermore, whenever (1.1) has an irreducible solution $\xi^{*}=p^{*}/q^{*}$, $((\eta_{\infty})^{{{2}}},p^{*},q^{*})$ is a solution of (2.5).
Proof.
First, it is true that ${\sqrt{\eta_{2}}}\leq\eta_{\infty}$. In fact, for any irreducible sequence $\{\xi^{(k)}=p^{(k)}/q^{(k)}\}$ satisfying $\eta_{\infty}^{(k)}:=\|\boldsymbol{f}-\xi^{(k)}(\boldsymbol{x})\|_{\infty}\rightarrow\eta_{\infty}$, $((\eta_{\infty}^{(k)})^{2},p^{(k)},q^{(k)})$ is feasible for (2.5), and thus ${\sqrt{\eta_{2}}}\leq\eta_{\infty}^{(k)}\rightarrow\eta_{\infty}$.
[…]
$|f_{j}\hat{q}(x_{j})-\hat{p}(x_{j})|^{{2}}=|\hat{p}(x_{j})|^{{2}}\leq 0\Longrightarrow\hat{p}(x_{j})=0,\quad\forall j=1,2,\dots,s.$
The proof is completed if $s=0$ (i.e., $\hat{q}(x_{j})\neq 0~\forall j=1,2,\dots,m$) because
$|f_{j}\hat{q}(x_{j})-\hat{p}(x_{j})|^{{2}}\leq{\eta_{2}}|\hat{q}(x_{j})|^{{2}}\Longrightarrow\left|f_{j}-\frac{\hat{p}(x_{j})}{\hat{q}(x_{j})}\right|\leq{\sqrt{\eta_{2}}}<\eta_{\infty},$
contradicting with the fact that $\eta_{\infty}$ is the infimum of (1.1).
We next consider $s\geq 1$. Choose a parameterized rational polynomial
$\frac{\hat{p}(x;\delta)}{\hat{q}(x;\delta)}=\frac{\hat{p}(x)+\delta\cdot p(x)}{\hat{q}(x)+\delta\cdot q(x)}\in\mathscr{R}_{(n_{1},n_{2})},~~\delta\in\mathbb{C},$
for some polynomials $p\in\mathbb{P}_{n_{1}}$ and $q\in\mathbb{P}_{n_{2}}$. We will finish the proof by showing that there is a pair $(p,q)$ and sufficiently small $\delta$ so that $\hat{q}(x_{j};\delta)\neq 0~\forall j=1,\dots,m,$ and
$\max_{1\leq j\leq m}\left|f_{j}-\frac{\hat{p}(x_{j};\delta)}{\hat{q}(x_{j};\delta)}\right|<\eta_{\infty},$ (2.6)
which contradicts with the fact that $\eta_{\infty}$ is the infimum of (1.1).
To this end, by Lemma 2.1, for $0<\epsilon:=\frac{\eta_{\infty}+\sqrt{\eta_{2}}}{2}<\eta_{\infty}$, there is a $\frac{p}{q}\in\mathscr{R}_{(n_{1},n_{2})}$ such that $\left|f_{j}-\frac{p(x_{j})}{q(x_{j})}\right|\leq\epsilon$ and $q(x_{j})\neq 0,~\forall j=1,\dots,s$. For any sufficiently small $\delta\neq 0$, we know that $\hat{q}(x_{j};\delta)\neq 0,~1\leq j\leq m$. Moreover, the parameterized $\frac{\hat{p}(x;\delta)}{\hat{q}(x;\delta)}$ with this pair $(p,q)$ satisfies
$\left|\frac{\hat{p}(x_{j};\delta)}{\hat{q}(x_{j};\delta)}-f_{j}\right|=\left|\frac{\hat{p}(x_{j})+\delta\cdot p(x_{j})}{\hat{q}(x_{j})+\delta\cdot q(x_{j})}-f_{j}\right|=\left|\frac{p(x_{j})}{q(x_{j})}-f_{j}\right|\leq\epsilon<\eta_{\infty},~\forall j=1,\dots,s.$
On the other hand, for any $j=s+1,\dots,m$, we have
$|f_{j}\hat{q}(x_{j})-\hat{p}(x_{j})|^{{2}}\leq{\eta_{2}}|\hat{q}(x_{j})|^{{2}}\Longrightarrow\left|f_{j}-\frac{\hat{p}(x_{j})}{\hat{q}(x_{j})}\right|\leq\sqrt{\eta_{2}}<\eta_{\infty},$
and
$\left|\frac{\hat{p}(x_{j})+\delta\cdot p(x_{j})}{\hat{q}(x_{j})+\delta\cdot q(x_{j})}-\frac{\hat{p}(x_{j})}{\hat{q}(x_{j})}\right|=|\delta|\cdot\left|\frac{p(x_{j})\hat{q}(x_{j})-q(x_{j})\hat{p}(x_{j})}{(\hat{q}(x_{j})+\delta\cdot q(x_{j}))\hat{q}(x_{j})}\right|\rightarrow 0,~~{\rm as}~~\delta\rightarrow 0.$
Consequently, for any sufficiently small $\delta$, it follows
$\left|\frac{\hat{p}(x_{j};\delta)}{\hat{q}(x_{j};\delta)}-f_{j}\right|$ $\leq\left|\frac{\hat{p}(x_{j};\delta)}{\hat{q}(x_{j};\delta)}-\frac{\hat{p}(x_{j})}{\hat{q}(x_{j})}\right|+\left|f_{j}-\frac{\hat{p}(x_{j})}{\hat{q}(x_{j})}\right|<\eta_{\infty},~\forall j=s+1,\dots,m.$
This leads to (2.6), and we have $\sqrt{\eta_{2}}=\eta_{\infty}$ by contradiction.
For the last part, according to $\sqrt{\eta_{2}}=\eta_{\infty}$ and Theorem 2.2, whenever (1.1) admits an irreducible solution $\xi^{*}=p^{*}/q^{*}$, the triplet $((\eta_{\infty})^{{2}},p^{*},q^{*})$ is a solution of (2.5). ∎
Theorem 2.3 reveals a profound connection, specifically $\eta_{2}=(\eta_{\infty})^{2}$, linking the rational minimax approximation problem (1.1) to its linearized counterpart in (2.5). Despite this significant relationship, there are instances where the attainability of infimum $\eta_{\infty}$ of (1.1) remains elusive and unattainable.
Example 2.1.
Consider $n_{1}=n_{2}=1$. Let $x_{j}=j~(1\leq j\leq 4)$, $f_{1}=0$ and $f_{j}=1$ for $j=2,3,4$. It is true that for any $\xi\in\mathscr{R}_{(1,1)}$, $\max_{1\leq j\leq 4}|f_{j}-\xi(x_{j})|>0$ because a nonconstant $\xi\in\mathscr{R}_{(1,1)}$ can only take each value once. However, the infimum $\eta_{\infty}=0$ since for the sequence $\xi^{(k)}(x)=\frac{x-1}{x-1+\frac{1}{k}}\in\mathscr{R}_{(1,1)}$, it holds that
$\max_{1\leq j\leq 4}|f_{j}-\xi^{(k)}(x_{j})|\rightarrow 0,~{\rm as}~k\rightarrow\infty.$
On the other hand, for (2.5), the infimum $\eta_{2}=0$ is attainable by p(x)=q(x)=x1p(x)=q(x)=x-1.
2.2 A dual problem
Even though we have transformed the original bi-level min-max problem (1.1) into (2.5), directly handling (2.5) is still hard. The idea in [54] is to develop the dual problem of (2.5), and then employ Lawson’s idea for the linear Chebyshev approximation problem for solving the dual problem. In particular, the dual function of (2.5) can be given by [54]
$d_{2}(\boldsymbol{w})$ $=\min_{\begin{subarray}{c}p\in\mathbb{P}_{n_{1}},~q\in\mathbb{P}_{n_{2}}\\ \sum_{j=1}^{m}w_{j}|q(x_{j})|^{2}=1\end{subarray}}\sum_{j=1}^{m}w_{j}|f_{j}q(x_{j})-p(x_{j})|^{2}$
$=\min_{\begin{subarray}{c}\boldsymbol{a}\in\mathbb{C}^{n_{1}+1},~\boldsymbol{b}\in\mathbb{C}^{n_{2}+1}\\ \|\sqrt{W}\Phi\boldsymbol{b}\|_{2}=1\end{subarray}}\left\|\sqrt{W}[-\Psi,F\Phi]\left[\begin{array}[]{c}\boldsymbol{a}\\ \boldsymbol{b}\end{array}\right]\right\|_{2}^{2},$ (2.9)
where $W={\rm diag}(\boldsymbol{w}),~F={\rm diag}(\boldsymbol{f})$. The following weak duality [54] has been proved:
$\forall\boldsymbol{w}\in{\cal S},~d_{2}(\boldsymbol{w})\leq(\eta_{\infty})^{2}\Longrightarrow\max_{\boldsymbol{w}\in{\cal S}}d_{2}(\boldsymbol{w})\leq(\eta_{\infty})^{2}.$ (2.10)
Moreover, by relying on Ruttan’s sufficient condition ([36, Theorem 2.1]; see also [25, Theorem 2] and [39, Theorem 3]), we can also have the following theoretical guarantee (see [54, Theorems 2.2, 4.3]) for solving the minimax approximant $\xi^{*}$.
Theorem 2.4.
Given $m\geq{n_{1}}+n_{2}+2$ distinct nodes ${\cal X}=\{x_{j}\}_{j=1}^{m}$ on $\Omega$, we have the weak duality (2.10). Let $\boldsymbol{w}^{*}\in{\cal S}$ be the solution to the dual problem
$\max_{\boldsymbol{w}\in{\cal S}}d_{2}(\boldsymbol{w}),$ (2.11)
and $(\boldsymbol{a}(\boldsymbol{w}^{*}),\boldsymbol{b}(\boldsymbol{w}^{*}))$ achieve the minimum $d_{2}(\boldsymbol{w}^{*})$ of (2.9). Then if the associated rational polynomial $\xi^{*}(x)=\frac{p^{*}(x)}{q^{*}(x)}=\frac{[\psi_{0}(x),\dots,\psi_{n_{1}}(x)]{\boldsymbol{a}(\boldsymbol{w}^{*})}}{[\phi_{0}(x),\dots,\phi_{n_{2}}(x)]{\boldsymbol{b}(\boldsymbol{w}^{*})}}$ is irreducible and
$\|\boldsymbol{f}-\xi^{*}(\boldsymbol{x})\|_{\infty}=\sqrt{d_{2}(\boldsymbol{w}^{*})},$ (2.12)
then $\xi^{*}$ is the minimax approximant of (1.1). When (2.12) holds, then we also have the following complementary slackness property:
$w_{j}^{*}\left(\|\boldsymbol{f}-\xi^{*}(\boldsymbol{x})\|_{\infty}-|f_{j}-\xi^{*}(x_{j})|\right)=0,\quad\forall j=1,2,\dots,m.$ (2.13)
We remark that the condition (2.12) implies the strong duality $\max_{\boldsymbol{w}\in{\cal S}}d_{2}(\boldsymbol{w})=(\eta_{\infty})^{2}$ [54, Theorem 4.3], which is equivalent to Ruttan’s sufficient condition for the original rational minimax problem (1.1). Therefore, under Ruttan’s sufficient condition, the complementary slackness property (2.13) necessarily holds at the maximizer $\boldsymbol{w}^{*}$ of (2.9). Furthermore, in the framework of dual programming (2.9), the accuracy of the associated approximation $\xi$ corresponding to the minimization (2.9) at an approximation $\boldsymbol{w}$ of $\boldsymbol{w}^{*}$, can be measured by the relative error
$\epsilon(\boldsymbol{w}):=\left|\frac{\sqrt{d_{2}(\boldsymbol{w})}-e(\xi)}{e(\xi)}\right|.$
This serves as a stopping rule for Lawson’s iteration (Algorithm 1).
2.3 Optimality for the dual objective function
To compute the dual function $d_{2}(\boldsymbol{w})$, a minimization problem (2.9) needs to be solved. The following proposition provides the optimality condition for this minimization.
Proposition 2.1.
([54, Proposition 3.1]) For $\boldsymbol{w}\in{\cal S}$, we have
  • (i)
    $\boldsymbol{c}(\boldsymbol{w})=\left[\begin{array}[]{c}\boldsymbol{a}(\boldsymbol{w})\\ \boldsymbol{b}(\boldsymbol{w})\end{array}\right]\in\mathbb{C}^{n_{1}+n_{2}+2}$ is a solution of (2.11) if and only if it is an eigenvector of the Hermitian positive semi-definite generalized eigenvalue problem $(A_{\boldsymbol{w}},B_{\boldsymbol{w}})$ and $d_{2}(\boldsymbol{w})$ is the smallest eigenvalue satisfying
    $A_{\boldsymbol{w}}\boldsymbol{c}(\boldsymbol{w})=d_{2}(\boldsymbol{w})B_{\boldsymbol{w}}\boldsymbol{c}(\boldsymbol{w})~\mbox{ and }~\boldsymbol{c}(\boldsymbol{w})^{\operatorname{H}}B_{\boldsymbol{w}}\boldsymbol{c}(\boldsymbol{w})=1,$ (2.14)
    where
    $A_{\boldsymbol{w}}:$ $=[-\Psi,F\Phi]^{\operatorname{H}}W[-\Psi,F\Phi]=\left[\begin{array}[]{cc}\Psi^{\operatorname{H}}W\Psi&-\Psi^{\operatorname{H}}FW\Phi\\ -\Phi^{\operatorname{H}}WF^{\operatorname{H}}\Psi&\Phi^{\operatorname{H}}F^{\operatorname{H}}WF\Phi\end{array}\right],$
    $B_{\boldsymbol{w}}:$ $=[0,\Phi]^{\operatorname{H}}W[0,\Phi]=\left[\begin{array}[]{cc}0&0\\ 0&\Phi^{\operatorname{H}}W\Phi\end{array}\right];$
  • (ii)
    the Hermitian matrix $H_{\boldsymbol{w}}:=A_{\boldsymbol{w}}-d_{2}(\boldsymbol{w})B_{\boldsymbol{w}}\succeq 0$, i.e., $H_{\boldsymbol{w}}$ is positive semi-definite;
  • (iii)
    let $W^{\frac{1}{2}}\Phi=Q_{q}R_{q}$ and $W^{\frac{1}{2}}\Psi=Q_{p}R_{p}$ be the thin QR factorizations where $Q_{q}\in\mathbb{C}^{m\times\widetilde{n}_{2}}$, $Q_{p}\in\mathbb{C}^{m\times\widetilde{n}_{1}}$, $R_{q}\in\mathbb{C}^{\widetilde{n}_{2}\times(n_{2}+1)}$, $R_{p}\in\mathbb{C}^{\widetilde{n}_{1}\times(n_{1}+1)}$ with $\widetilde{n}_{1}={\rm rank}(W^{\frac{1}{2}}\Psi)$ and $\widetilde{n}_{2}={\rm rank}(W^{\frac{1}{2}}\Phi)$. Then $(d_{2}(\boldsymbol{w}),R_{q}\boldsymbol{b}(\boldsymbol{w}))$ is an eigenpair associated with the smallest eigenvalue of the Hermitian positive semi-definite matrix $S_{F}-S_{qp}S_{qp}^{\operatorname{H}}\in\mathbb{C}^{\widetilde{n}_{2}\times\widetilde{n}_{2}}$ satisfying
    $(S_{F}-S_{qp}S_{qp}^{\operatorname{H}})R_{q}\boldsymbol{b}(\boldsymbol{w})$ $=d_{2}(\boldsymbol{w})R_{q}\boldsymbol{b}(\boldsymbol{w}),~~\|R_{q}\boldsymbol{b}(\boldsymbol{w})\|_{2}=1,$ (2.15a)
    $R_{p}\boldsymbol{a}(\boldsymbol{w})$ $=S_{qp}^{\operatorname{H}}R_{q}\boldsymbol{b}(\boldsymbol{w}),$ (2.15b)
    where $S_{F}=Q_{q}^{\operatorname{H}}|F|^{2}Q_{q}\in\mathbb{C}^{\widetilde{n}_{2}\times\widetilde{n}_{2}},~S_{qp}=Q_{q}^{\operatorname{H}}F^{\operatorname{H}}Q_{p}\in\mathbb{C}^{\widetilde{n}_{2}\times\widetilde{n}_{1}}.$ Moreover, letting $Q_{p},Q_{p}^{\perp}]\in\mathbb{C}^{m\times m}$ be unitary, then $(\sqrt{d_{2}(\boldsymbol{w})},R_{q}\boldsymbol{b}(\boldsymbol{w}))$ is the right singular pair associated with the smallest singular value of both $(Q_{p}^{\perp})^{\operatorname{H}}FQ_{q}\in\mathbb{C}^{(m-\widetilde{n}_{1}-1)\times(\widetilde{n}_{2}+1)}$ and $(I_{m}-Q_{p}Q_{p}^{\operatorname{H}})FQ_{q}\in\mathbb{C}^{m\times(\widetilde{n}_{2}+1)}$.
As remarked in [54], the normalized condition $\boldsymbol{c}(\boldsymbol{w})^{\operatorname{H}}B_{\boldsymbol{w}}\boldsymbol{c}(\boldsymbol{w})=1$ in (2.14) can always be fulfilled whenever $q\not\equiv 0$ and there are at least $n_{2}+1$ positive elements in $w_{j}$. In fact, $0=\boldsymbol{c}(\boldsymbol{w})^{\operatorname{H}}B_{\boldsymbol{w}}\boldsymbol{c}(\boldsymbol{w})=\sum_{j=1}^{m}w_{j}|q(x_{j})|^{2}\Longrightarrow w_{j}q(x_{j})=0,~{\forall j=1,2,\dots,m};$ thus, any node $x_{j}$ with $w_{j}>0$ is a zero of $q$, and if $\boldsymbol{w}$ has at least $n_{2}+1$ positive elements, it implies that $q\equiv 0$. Thus, in the following discussion, we assume that any iterate $\boldsymbol{w}^{(k)}$ has at least $n_{2}+1$ positive elements.
We remark further that, by rewriting the first $n_{1}+1$ rows and the last $n_{2}+1$ rows of the optimality condition $A_{\boldsymbol{w}}\boldsymbol{c}(\boldsymbol{w})=d_{2}(\boldsymbol{w})B_{\boldsymbol{w}}\boldsymbol{c}(\boldsymbol{w})$ in (2.14), we have
Corollary 2.1.
Let $\boldsymbol{p}=\Psi\boldsymbol{a}(\boldsymbol{w})$ and $\boldsymbol{q}=\Phi\boldsymbol{b}(\boldsymbol{w})$ be from the solution of (2.9) with the weight vector $\boldsymbol{w}$. Then
$F\boldsymbol{q}-\boldsymbol{p}\perp_{\boldsymbol{w}}{\rm span}(\Psi),~~F^{\operatorname{H}}(F\boldsymbol{q}-\boldsymbol{p})-d_{2}(\boldsymbol{w})\boldsymbol{q}\perp_{\boldsymbol{w}}{\rm span}(\Phi).$ (2.16)
Besides the optimality in solving the minimization (2.9) for the minimum $d_{2}(\boldsymbol{w})$, we can further obtain the gradient of $d_{2}(\boldsymbol{w})$ with respect to the dual variable $\boldsymbol{w}$.
Proposition $2.$
([54, Proposition 5.1]) For $\boldsymbol{w}>0$, let $d_{2}(\boldsymbol{w})$ be the smallest eigenvalue of the Hermitian positive semi-definite generalized eigenvalue problem (2.14), and $\boldsymbol{c}(\boldsymbol{w})=\left[\begin{array}[]{c}\boldsymbol{a}(\boldsymbol{w})\\ \boldsymbol{b}(\boldsymbol{w})\end{array}\right]\in\mathbb{C}^{n_{1}+n_{2}+2}$ be the associated eigenvector. Denote
$\boldsymbol{p}=[p_{1},\dots,p_{m}]^{\operatorname{T}}=\Psi\boldsymbol{a}(\boldsymbol{w})\in\mathbb{C}^{m}~{\rm and}~\boldsymbol{q}=[q_{1},\dots,q_{m}]^{\operatorname{T}}=\Phi\boldsymbol{b}(\boldsymbol{w})\in\mathbb{C}^{m}.$
If $d_{2}(\boldsymbol{w})$ is a simple eigenvalue, then $d_{2}(\boldsymbol{w})$ is differentiable with respect to $\boldsymbol{w}$ and its gradient is
$\nabla d_{2}(\boldsymbol{w})=\left[\begin{array}[]{c}|f_{1}q_{1}-p_{1}|^{2}-d_{2}(\boldsymbol{w})|q_{1}|^{2}\\ |f_{2}q_{2}-p_{2}|^{2}-d_{2}(\boldsymbol{w})|q_{2}|^{2}\\ \vdots\\ |f_{m}q_{m}-p_{m}|^{2}-d_{2}(\boldsymbol{w})|q_{m}|^{2}\end{array}\right]=:|F\boldsymbol{q}-\boldsymbol{p}|^{2}-d_{2}(\boldsymbol{w})|\boldsymbol{q}|^{2}\in\mathbb{R}^{m}.$ (2.17)
2.4 The d-Lawson iteration
Within the framework of the dual problem, it has been claimed [54] that Lawson’s iteration is a method for solving the dual problem (2.11). For the rational minimax approximation problem, Lawson’s iteration [54] is implemented as in Algorithm 1. Numerical results have been reported in [54], indicating that d-Lawson [^1] The MATLAB code of d-Lawson is available at https://ww2.mathworks.cn/matlabcentral/fileexchange/167176-d-lawson-method. generally converges monotonically with respect to the dual function value $d_{2}(\boldsymbol{w})$. In our following discussion, we assume $\epsilon_{w}=0$ in Step 1 of Algorithm 1.
Algorithm 1 A rational Lawson’s iteration (d-Lawson) [54] for (1.2)
0: Given samples $\{(x_{j},f_{j})\}_{j=1}^{m}$ and $0\leq n_{1}+n_{2}+2\leq m$ with $x_{j}\in\Omega$, a relative tolerance for the strong duality $\epsilon_{r}>0$, the maximum number $k_{\rm maxit}$ of iterations;
1: (Initialization) Let $k=0$; choose $0<\boldsymbol{w}^{(0)}\in{\cal S}$ and a tolerance $\epsilon_{w}$ for the weights;
2: (Filtering) Remove nodes $x_{i}$ with $w_{i}^{(k)}<\epsilon_{w}$;
3: Compute $d_{2}(\boldsymbol{w}^{(k)})$ and the associated vector $\xi^{(k)}(\boldsymbol{x})={[\Psi\boldsymbol{a}(\boldsymbol{w}^{(k)})]./[\Phi\boldsymbol{b}(\boldsymbol{w}^{(k)})]}$ according to Proposition 2.1;
4: (Stopping rule) Stop either if $k\geq k_{\rm maxit}$ or
$\epsilon(\boldsymbol{w}^{(k)}):=\left|\frac{\sqrt{d_{2}(\boldsymbol{w}^{(k)})}-e(\xi^{{(k)}})}{e(\xi^{{(k)}})}\right|<\epsilon_{r},~~{\rm where}~~e(\xi^{{(k)}})=\|\boldsymbol{f}-\xi^{(k)}(\boldsymbol{x})\|_{\infty};$
5: (Updating weights) Update the weight vector $\boldsymbol{w}^{(k+1)}$ according to
$w_{j}^{(k+1)}=\frac{w_{j}^{(k)}\left|f_{j}-\xi^{(k)}(x_{j})\right|^{\beta}}{\sum_{i}w_{i}^{(k)}\left|f_{i}-\xi^{(k)}(x_{i})\right|^{\beta}},~~\forall j,$ (2.18)
with the Lawson exponent $\beta>0$, and go to Step 2 with $k=k+1$.
Remark 2.1.
It is interesting to point out that d-Lawson in Algorithm 1 naturally reduces to the classical Lawson’s iteration [27] for the linear (polynomial) minimax approximation problem when $n_{2}=0$, which corresponds to $q\equiv 1$ and $\boldsymbol{q}^{\operatorname{H}}W\boldsymbol{q}\equiv 1$. This observation unifies our subsequent convergence analysis.
Remark 2.2.
In practice, the implementation of Step 3 should handle the stability and accuracy for computing $d_{2}(\boldsymbol{w}^{(k)})$ and the associated vector $\xi^{(k)}(\boldsymbol{x})=\boldsymbol{p}^{(k)}./\boldsymbol{q}^{(k)}$ where $\boldsymbol{p}^{(k)}=\Psi\boldsymbol{a}(\boldsymbol{w}^{(k)})$ and $\boldsymbol{q}^{(k)}=\Phi\boldsymbol{b}(\boldsymbol{w}^{(k)})$. In [54], the Vandermonde with Arnoldi process [4, 24, 53] is employed for this step.
It is noticed that Lawson’s updating (2.18) relies on the error vector $\boldsymbol{f}-\xi^{(k)}(\boldsymbol{x})$ to update the weights. In this procedure, each entry of the denominator vector $\boldsymbol{q}^{(k)}$ is assumed to be nonzero; this is generally the case in practice, and in our following discussion, we make the following two assumptions:
  • (A1)
    $q_{j}^{(k)}=\boldsymbol{e}_{j}^{\operatorname{T}}\boldsymbol{q}^{(k)}\neq 0$ for any $j=1,2,\dots,m$ and any $k\geq 0$;
  • (A2)
    the cardinality $|{\cal I}_{\boldsymbol{w}^{(k)}}|\geq\max\{n_{1}+1,n_{2}+1\}$, where
    ${\cal I}_{\boldsymbol{w}^{(k)}}:=\{j|w_{j}^{(k)}\neq 0,~{1\leq j\leq m}\}.$
Proposition 2.3.
If $d_{2}(\boldsymbol{w}^{(k)})>0$, then
${\cal I}_{\boldsymbol{w}^{(k+1)}}=\{j|w_{j}^{(k)}r_{j}^{(k)}\neq 0,~{1\leq j\leq m}\},~~{\rm where}~~r_{j}^{(k)}=|f(x_{j})-\xi^{(k)}(x_{j})|.$
Proof.
By (A1), the condition $d_{2}(\boldsymbol{w}^{(k)})>0$ implies $\sum_{i}w_{i}^{(k)}\left|f_{i}-\xi^{(k)}(x_{j})\right|^{\beta}>0$. Thus, according to Lawson’s updating rule (2.18), the conclusion follows due to the fact that $w_{j}^{(k+1)}=0$ is equivalent to $w_{j}^{(k)}r_{j}^{(k)}=0$. ∎
2.5 The roadmap of the convergence analysis
Before proceeding the detailed convergence analysis for d-Lawson, we present a roadmap of our analysis in Fig. 2.1 which can help the reader to follow the overly technical presentation and make the main novelty of our theoretical analysis clearer.
[…]
Figure 2.1: The roadmap of the convergence analysis for Lawson’s iteration, where $d_{2}^{(k)}=d_{2}(\boldsymbol{w}^{(k)})$
From this roadmap in Fig. 2.1, we remark that the main idea is to establish the monotonic property $d_{2}({\boldsymbol{w}^{(k+1)}})\geq d_{2}({\boldsymbol{w}^{(k)}})$ with respect to the dual objective function. In this framework, the dual problem (2.11) plays the most important role, whose solution corresponds to the minimax approximant of the original (1.1) under Ruttan’s sufficient condition (see Theorem 2.4). The remarkable point is that d-Lawson serves as a simple yet efficient monotonically ascend method for the dual (2.11).
Based on the roadmap in Fig. 2.1, we also make a remark on the AAA-Lawson iteration [14]. Comparing with d-Lawson, the main differences between these two methods are: (1) AAA-Lawson uses barycentric representations for $\xi^{(k)}$ and, (2) for a given $\boldsymbol{w}^{(k)}\in{\cal S}$, they compute the associated $\xi^{(k)}$ in different ways. We remark that barycentric representations can be very useful for numerical stability, but mathematically (i.e., in exact arithmetic) the dual framework does not rely on a particular representation for $\xi^{(k)}$; in fact, barycentric representations can also be adopted in representing (2.1) and the d-Lawson version with barycentric representations is one of our ongoing works. The crucial difference between AAA-Lawson and d-Lawson is the way of computing the new rational approximant $\xi^{(k)}$ at a given $\boldsymbol{w}^{(k)}$; in particular, AAA-Lawson defines $\xi^{(k)}$ as
$\xi_{\tiny{\rm AAA-Lawson}}^{(k)}=\frac{p^{(k)}}{q^{(k)}},~(p^{(k)},q^{(k)})=\operatorname*{argmin}_{\begin{subarray}{c}p\in\mathbb{P}_{n_{1}},q\in\mathbb{P}_{n_{2}}\\ \|\boldsymbol{a}\|_{2}^{2}+\|\boldsymbol{b}\|_{2}^{2}=1\end{subarray}}\sum_{j=1}^{m}w^{(k)}_{j}|f_{j}q(x_{j})-p(x_{j})|^{2},$ (2.19)
where $\boldsymbol{a}$ and $\boldsymbol{b}$ are coefficient vectors of $p^{(k)}$ and $q^{(k)}$, respectively, in barycentric representations, while d-Lawson defines $\xi^{(k)}$ as
$\xi^{(k)}$ $=\frac{p^{(k)}}{q^{(k)}},~(p^{(k)},q^{(k)})=\operatorname*{argmin}_{\begin{subarray}{c}p\in\mathbb{P}_{n_{1}},~q\in\mathbb{P}_{n_{2}}\\ \sum_{j=1}^{m}w^{(k)}_{j}|q(x_{j})|^{2}=1\end{subarray}}\sum_{j=1}^{m}w^{(k)}_{j}|f_{j}q(x_{j})-p(x_{j})|^{2}.$
It appears that $\xi_{\tiny{\rm AAA-Lawson}}^{(k)}$ is a seeming extension of the linear minimax approximation problem to the rational case which lacks an underlying dual problem. When returning to the linear case with $n_{2}=0$, both reduce to Lawson’s original form (see Remark 2.1), but for the rational case $n_{2}>0$, $\xi^{(k)}$ is the more powerful formulation and generalization because it is derived from the dual framework. We refer the reader to [54] for a comparison of the numerical performance of these two methods in terms of accuracy and robustness.
3 A lower bound of the dual objective function value
We first relate Lawson’s iteration with the gradient ascent direction. Indeed, when the current $0<\boldsymbol{w}\in{\cal S}$, $q_{j}=\boldsymbol{q}^{\operatorname{T}}\boldsymbol{e}_{j}\neq 0~{\forall j=1,2,\dots,m}$ and $d_{2}(\boldsymbol{w})$ is the simple eigenvalue of $(A_{\boldsymbol{w}},B_{\boldsymbol{w}})$, then we know that the direction
$\boldsymbol{g}(\boldsymbol{w}):={\rm diag}\left(\frac{w_{1}}{|q_{1}|^{2}},\ldots,\frac{w_{m}}{|q_{m}|^{2}}\right)\nabla d_{2}(\boldsymbol{w})\in\mathbb{R}^{m}$
is an ascent direction for the dual function $d_{2}(\boldsymbol{w})$ because by (2.17)
$\boldsymbol{g}(\boldsymbol{w})^{\operatorname{T}}\nabla d_{2}(\boldsymbol{w})=\nabla d_{2}(\boldsymbol{w})^{\operatorname{T}}{\rm diag}\left(\frac{w_{1}}{|q_{1}|^{2}},\ldots,\frac{w_{m}}{|q_{m}|^{2}}\right)\nabla d_{2}(\boldsymbol{w})>0.$
Note that using this direction, the update with step size $\mu=\frac{1}{d_{2}(\boldsymbol{w})}>0$ gives
$\widetilde{\boldsymbol{w}}:=\boldsymbol{w}+\mu\boldsymbol{g}={\rm diag}(\boldsymbol{w})|\boldsymbol{f}-\xi(\boldsymbol{x})|^{2}/d_{2}(\boldsymbol{w})$
which, after a scaling $\widetilde{\boldsymbol{w}}\leftarrow\frac{\widetilde{\boldsymbol{w}}}{\boldsymbol{e}^{\operatorname{T}}\widetilde{\boldsymbol{w}}}\in{\cal S}$, implies the iteration is the same as Lawson’s iteration (2.18) with $\beta=2$. Note that the scaling $\widetilde{\boldsymbol{w}}\leftarrow\frac{\widetilde{\boldsymbol{w}}}{\boldsymbol{e}^{\operatorname{T}}\widetilde{\boldsymbol{w}}}$ can be viewed as a certain projection onto $\cal S$}. From this point of view, we can say that a Lawson’s iteration in Algorithm 1 with the Lawson exponent $\beta=2$ is just an ascent gradient step with a specific step-size followed by a certain projection onto $\cal S$}.
To more clearly see the relation between the two consecutive objective values $d_{2}(\boldsymbol{w}^{(k)})$ and $d_{2}(\boldsymbol{w}^{(k+1)})$, we introduce the $\boldsymbol{w}$-inner (positive semidefinite) product defined by $\langle\boldsymbol{y},\boldsymbol{z}\rangle_{\boldsymbol{w}}=\boldsymbol{y}^{\operatorname{H}}W\boldsymbol{z}$ and $\|\boldsymbol{y}\|_{\boldsymbol{w}}=\sqrt{\boldsymbol{y}^{\operatorname{H}}W\boldsymbol{y}}$, where $W={\rm diag}(\boldsymbol{w})$ and $\boldsymbol{w}\geq 0$. The following lemma, which is a generalized result of the standard least-squares problem, plays an important role in establishing the convergence of Lawson’s iteration [27] for the linear minimax approximation problem [33, Lemma 13-12].
Lemma 3.1.
Given $0\leq\boldsymbol{w}\in\mathbb{R}^{m}$ satisfying $|{\cal I}_{\boldsymbol{w}}|\geq n$ with ${\cal I}_{\boldsymbol{w}}=\{j|w_{j}>0\}$, a matrix $A\in\mathbb{C}^{m\times n}$ with ${\rm rank}(A({\cal I}_{\boldsymbol{w}},:))=n$ and $\boldsymbol{z}\in\mathbb{C}^{m}$, let $\boldsymbol{x}_{*}$ be the solution to the least-squares problem
$\min_{\boldsymbol{x}\in{\rm span}(A)}\|\boldsymbol{x}-\boldsymbol{z}\|_{\boldsymbol{w}}.$
Then we have
$\frac{\boldsymbol{x}_{*}-\boldsymbol{z}}{\|\boldsymbol{x}_{*}-\boldsymbol{z}\|_{\boldsymbol{w}}}\in\arg\max_{\|\boldsymbol{y}\|_{\boldsymbol{w}}=1,~\boldsymbol{y}^{\operatorname{H}}WA=0}|\langle\boldsymbol{z},\boldsymbol{y}\rangle_{\boldsymbol{w}}|.$
Proof.
Let $W={\rm diag}(\boldsymbol{w})$. We first consider the case $\boldsymbol{w}>0$.
Denote by $P\in\mathbb{C}^{m\times(m-n)}$ the $\boldsymbol{w}$-orthogonal basis for the complement [26, Chapter 1.6.3] of ${\rm span}(A)$ satisfying ${\rm span}(P)\oplus{\rm span}(A)=\mathbb{C}^{m}$, $P^{\operatorname{H}}WP=I_{m-n}$ and $P^{\operatorname{H}}WA=0$. Note that $PP^{\operatorname{H}}W$ and $I-PP^{\operatorname{H}}W$ are projections onto ${\rm span}(P)$ and ${\rm span}(A)$, respectively. For the least-squares problem, we know that $\boldsymbol{x}_{*}$ is the solution if and only if
$(\boldsymbol{x}_{*}-\boldsymbol{z})^{\operatorname{H}}WA=0,~~(i.e.,~\boldsymbol{x}_{*}-\boldsymbol{z}=PP^{\operatorname{H}}W\boldsymbol{z}).$
Note $\|\boldsymbol{x}_{*}-\boldsymbol{z}\|_{\boldsymbol{w}}=\|PP^{\operatorname{H}}W\boldsymbol{z}\|_{\boldsymbol{w}}=\|P^{\operatorname{H}}W\boldsymbol{z}\|_{2}$. Also, the constraint $\|\boldsymbol{y}\|_{\boldsymbol{w}}=1,~\boldsymbol{y}^{\operatorname{H}}WA=0$ can be parameterized as $\boldsymbol{y}=P\boldsymbol{t}$ for $\boldsymbol{t}\in\mathbb{C}^{m-n}$ with $\|\boldsymbol{t}\|_{2}=1$. Since
$|\langle\boldsymbol{z},\boldsymbol{y}\rangle_{\boldsymbol{w}}|=|\boldsymbol{z}^{\operatorname{H}}WP\boldsymbol{t}|\leq\|P^{\operatorname{H}}W\boldsymbol{z}\|_{2},$
and the equality holds if $\boldsymbol{t}=P^{\operatorname{H}}W\boldsymbol{z}/\|P^{\operatorname{H}}W\boldsymbol{z}\|_{2}$, it implies that
$\boldsymbol{y}=\frac{PP^{\operatorname{H}}W\boldsymbol{z}}{\|P^{\operatorname{H}}W\boldsymbol{z}\|_{2}}=\frac{\boldsymbol{x}_{*}-\boldsymbol{z}}{\|\boldsymbol{x}_{*}-\boldsymbol{z}\|_{\boldsymbol{w}}}$
is the maximizer and the conclusion follows.
For the general case $\boldsymbol{w}\geq 0$, assume ${\cal I}_{\boldsymbol{w}}=\{1,2,\dots,t\}$, and partition accordingly $\boldsymbol{x}=[\boldsymbol{x}_{1}^{\operatorname{T}},\boldsymbol{x}_{2}^{\operatorname{T}}]^{\operatorname{T}}$, $\boldsymbol{z}=[\boldsymbol{z}_{1}^{\operatorname{T}},\boldsymbol{z}_{2}^{\operatorname{T}}]^{\operatorname{T}}$, $\boldsymbol{w}=[\boldsymbol{w}_{1}^{\operatorname{T}},\boldsymbol{w}_{2}^{\operatorname{T}}]^{\operatorname{T}}$, $\boldsymbol{y}=[\boldsymbol{y}_{1}^{\operatorname{T}},\boldsymbol{y}_{2}^{\operatorname{T}}]^{\operatorname{T}}$, and $A=[A_{1}^{\operatorname{T}},A_{2}^{\operatorname{T}}]^{\operatorname{T}}$ with $A_{1}\in\mathbb{C}^{t\times n}$ and ${\rm rank}(A_{1})=n$. Note
$\min_{\boldsymbol{x}\in{\rm span}(A)}\|\boldsymbol{x}-\boldsymbol{z}\|_{\boldsymbol{w}}=\min_{\boldsymbol{x}_{1}\in{\rm span}(A_{1})}\|\boldsymbol{x}_{1}-\boldsymbol{z}_{1}\|_{\boldsymbol{w}_{1}},$
and the solution $\boldsymbol{x}_{*}$ is unique by assumptions. Indeed, $\boldsymbol{x}_{*}$ satisfies $(\boldsymbol{x}_{*}-\boldsymbol{z})^{\operatorname{H}}WA=0.$
On the other hand,
$\max_{\|\boldsymbol{y}\|_{\boldsymbol{w}}=1,~\boldsymbol{y}^{\operatorname{H}}WA=0}|\langle\boldsymbol{z},\boldsymbol{y}\rangle_{\boldsymbol{w}}|=\max_{\|\boldsymbol{y}_{1}\|_{\boldsymbol{w}_{1}}=1,~\boldsymbol{y}_{1}^{\operatorname{H}}W_{1}A_{1}=0}|\langle\boldsymbol{z}_{1},\boldsymbol{y}_{1}\rangle_{\boldsymbol{w}_{1}}|$
and by the proof of the first case $\boldsymbol{w}>0$, we know that any maximizer $\boldsymbol{y}=[\boldsymbol{y}_{1}^{\operatorname{T}},\boldsymbol{y}_{2}^{\operatorname{T}}]^{\operatorname{T}}$ takes $\boldsymbol{y}_{1}=\frac{\boldsymbol{x}{{}_{1}}_{*}-\boldsymbol{z}_{1}}{\|\boldsymbol{x}{{}_{1}}_{*}-\boldsymbol{z}_{1}\|_{\boldsymbol{w}_{1}}}$ and $\forall\boldsymbol{y}_{2}\in\mathbb{C}^{m-t}$. Thus, the conclusion follows. ∎
Lemma 3.2.
Let $(\boldsymbol{a}(\boldsymbol{w}^{(k)}),\boldsymbol{b}(\boldsymbol{w}^{(k)}))$ be the solution and $d_{2}(\boldsymbol{w}^{(k)})$ be the minimum of (2.9) with the weight $\boldsymbol{w}^{(k)}\in{\cal S}$. Denote $\boldsymbol{p}^{(k)}=\Psi\boldsymbol{a}(\boldsymbol{w}^{(k)}),~\boldsymbol{q}^{(k)}=\Phi\boldsymbol{b}(\boldsymbol{w}^{(k)})$ and $\xi^{(k)}(\boldsymbol{x})=\boldsymbol{p}^{(k)}./\boldsymbol{q}^{(k)}$. If $d_{2}(\boldsymbol{w}^{(k)})>0$, then
$(F\boldsymbol{q}^{(k)}-\boldsymbol{p}^{(k)})./|\boldsymbol{f}-\xi^{(k)}(\boldsymbol{x})|^{{\beta}}\perp_{\boldsymbol{w}^{(k+1)}}{\rm span}(\Psi).$ (3.1)
Here we set $\frac{0}{0}=0$ for convenience.
Proof.
[…]
Notice that $d_{2}(\boldsymbol{w})>0$ implies
${\cal I}_{\widetilde{\boldsymbol{w}}}=\{j|\widetilde{w}_{j}\neq 0,~{1\leq j\leq m}\}=\{j|w_{j}r_{j}\neq 0,~{1\leq j\leq m}\}$
by Proposition 2.3. Also, according to (2.18), $\widetilde{w}_{j}=w_{j}|f_{j}-\xi_{j}|^{\beta}/\gamma_{\beta},$ where
$\gamma_{\beta}:=\sum_{j=1}^{m}w_{j}\left|f_{j}-\xi(x_{j})\right|^{\beta}\left(=\left\||\boldsymbol{f}-\xi(\boldsymbol{x})|^{\frac{\beta}{2}}\right\|_{\boldsymbol{w}}^{2}\right)>0.$
[…]
$\min_{\boldsymbol{a}\in\mathbb{C}^{n_{1}+1}}\left\|F\widetilde{\boldsymbol{q}}-\Psi\boldsymbol{a}\right\|_{\widetilde{\boldsymbol{w}}}=\min_{\boldsymbol{z}\in{\rm span}(\Psi)}\left\|F\widetilde{\boldsymbol{q}}-\boldsymbol{z}\right\|_{\widetilde{\boldsymbol{w}}}$;
thus we have the optimality $(F\widetilde{\boldsymbol{q}}-\widetilde{\boldsymbol{p}})^{\operatorname{H}}\widetilde{W}\Psi=0$ (see (2.16)); analogously, for the previous iteration, we have $(F\boldsymbol{q}-\boldsymbol{p})^{\operatorname{H}}W\Psi=0$ implying that $\forall\boldsymbol{t}\in{\rm span}(\Psi)$,
$0=\langle\boldsymbol{t},F\boldsymbol{q}-\boldsymbol{p}\rangle_{\boldsymbol{w}}$ $=\sum_{j=1}^{m}w_{j}\bar{t}_{j}q_{j}(f_{j}-\xi_{j})=\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}\bar{t}_{j}q_{j}(f_{j}-\xi_{j})$
${=\gamma_{\beta}\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}\widetilde{w}_{j}\bar{t}_{j}q_{j}(f_{j}-\xi_{j})|f_{j}-\xi_{j}|^{-\beta}}$
$=\gamma_{\beta}\sum_{j=1}^{m}\widetilde{w}_{j}\bar{t}_{j}q_{j}(f_{j}-\xi_{j})|f_{j}-\xi_{j}|^{-\beta}\quad\quad\mbox{(by ~$\widetilde{w}_{j}=0~\forall j\not\in{\cal I}_{\widetilde{\boldsymbol{w}}}$)}$
$=\gamma_{\beta}\langle\boldsymbol{t},(F\boldsymbol{q}-\boldsymbol{p})./|\boldsymbol{f}-\xi(\boldsymbol{x})|^{{\beta}}\rangle_{\widetilde{\boldsymbol{w}}}$.
Thus, $(F\boldsymbol{q}-\boldsymbol{p})./|\boldsymbol{f}-\xi(\boldsymbol{x})|^{{\beta}}\perp_{\widetilde{\boldsymbol{w}}}{\rm span}(\Psi)$. ∎
We next provide a lower bound for the dual objective function value $d_{2}(\boldsymbol{w}^{(k+1)})$. This lower bound plays a crucial role in finding the near-optimal $\beta$ for Lawson’s iteration in the linear minimax approximation problem, and also is a key to establish monotonic convergence $d_{2}(\boldsymbol{w}^{(k+1)})\geq d_{2}(\boldsymbol{w}^{(k)})$ in the rational minimax approximation problem.
Theorem 3.1.
Let $(\boldsymbol{a}(\boldsymbol{w}^{(k)}),\boldsymbol{b}(\boldsymbol{w}^{(k)}))$ be the solution and $d_{2}(\boldsymbol{w}^{(k)})$ be the minimum of (2.9) with the weight $\boldsymbol{w}^{(k)}\in{\cal S}$. Denote $\boldsymbol{p}^{(k)}=\Psi\boldsymbol{a}(\boldsymbol{w}^{(k)}),~\boldsymbol{q}^{(k)}=\Phi\boldsymbol{b}(\boldsymbol{w}^{(k)})$ and $\xi^{(k)}(\boldsymbol{x})=\boldsymbol{p}^{(k)}./\boldsymbol{q}^{(k)}$. Then
$\sqrt{d_{2}(\boldsymbol{w}^{(k+1)})}\geq d_{2}(\boldsymbol{w}^{(k)})\cdot\chi^{(k)}(\beta),$ (3.2)
where $\chi^{(k)}(\beta):=\frac{\left|(\boldsymbol{q}^{(k+1)})^{\operatorname{H}}W^{(k)}\boldsymbol{q}^{(k)}\right|}{\gamma_{\beta}~\zeta_{\beta}}$, $W^{(k)}={\rm diag}(\boldsymbol{w}^{(k)})$,
$\gamma_{\beta}=\left\||\boldsymbol{f}-\xi^{(k)}(\boldsymbol{x})|^{\frac{\beta}{2}}\right\|_{\boldsymbol{w}^{(k)}}^{2}:=\sum_{j=1}^{m}w_{j}^{(k)}\left|f_{j}-\xi^{(k)}(x_{j})\right|^{\beta},~\zeta_{\beta}=\left\|\frac{(\boldsymbol{f}\boldsymbol{q}^{(k)}-\boldsymbol{p}^{(k)}).}{\left|\boldsymbol{f}-\xi^{(k)}(\boldsymbol{x})\right|^{{\beta}}}\right\|_{\boldsymbol{w}^{(k+1)}},$
and we set $\frac{0}{0}=0$ for convenience.
Proof.
Following the notation in the proof of Lemma 3.2, we drop the superscript $k$ in each quantity related with the $k$th iteration. The result of (3.2) is trivial if $d_{2}(\boldsymbol{w})=0$, and we assume $d_{2}(\boldsymbol{w})>0$, which, by Proposition 2.3, implies ${\cal I}_{\widetilde{\boldsymbol{w}}}=\{j|\widetilde{w}_{j}\neq 0,~{1\leq j\leq m}\}=\{j|w_{j}r_{j}\neq 0,~{1\leq j\leq m}\}.$
By assumptions (A1), (A2), $\widetilde{w}_{j}=w_{j}|f_{j}-\xi_{j}|^{\beta}/\gamma_{\beta}$, and the fact that any $n_{1}+1$ rows of $\Psi$ are linearly independent, we can apply Lemmas 3.1 and 3.2 to get
[…]
$=\frac{\sqrt{d_{2}(\widetilde{\boldsymbol{w}})}}{\zeta_{\beta}}~\left|\sum_{j=1}^{m}\widetilde{w}_{j}\overline{f_{j}\widetilde{q_{j}}}(f_{j}q_{j}-p_{j})/|f_{j}-\xi_{j}|^{\beta}\right|$
$=\frac{\sqrt{d_{2}(\widetilde{\boldsymbol{w}})}}{\gamma_{\beta}\zeta_{\beta}}~\left|\sum_{j=1}^{m}w_{j}\overline{f_{j}\widetilde{q_{j}}}(f_{j}q_{j}-p_{j})\right|$
$=\frac{\sqrt{d_{2}(\widetilde{\boldsymbol{w}})}}{\gamma_{\beta}\zeta_{\beta}}~\left|\langle F\widetilde{\boldsymbol{q}},F\boldsymbol{q}-\boldsymbol{p}\rangle_{\boldsymbol{w}}\right|=\frac{\sqrt{d_{2}(\widetilde{\boldsymbol{w}})}}{\gamma_{\beta}\zeta_{\beta}}{d_{2}(\boldsymbol{w})\left|\widetilde{\boldsymbol{q}}^{\operatorname{H}}W\boldsymbol{q}\right|},$
where the last equality $\langle F\widetilde{\boldsymbol{q}},F\boldsymbol{q}-\boldsymbol{p}\rangle_{\boldsymbol{w}}=d_{2}(\boldsymbol{w})\widetilde{\boldsymbol{q}}^{\operatorname{H}}W\boldsymbol{q}$ is due to the second optimality formula in (2.16) for the pair $(\boldsymbol{p},\boldsymbol{q})$:
$F^{\operatorname{H}}(F\boldsymbol{q}-\boldsymbol{p})-d_{2}(\boldsymbol{w})\boldsymbol{q}\perp_{\boldsymbol{w}}{\rm span}(\Phi)$ $\Longrightarrow\widetilde{\boldsymbol{b}}^{\operatorname{H}}\Phi^{\operatorname{H}}W(F^{\operatorname{H}}(F\boldsymbol{q}-\boldsymbol{p})-d_{2}(\boldsymbol{w})\boldsymbol{q})=0$
$\Longrightarrow\langle F\widetilde{\boldsymbol{q}},F\boldsymbol{q}-\boldsymbol{p}\rangle_{\boldsymbol{w}}=d_{2}(\boldsymbol{w})\widetilde{\boldsymbol{q}}^{\operatorname{H}}W\boldsymbol{q}.$
The proof is complete. ∎
4 $\beta=1$ is near-optimal for the linear minimax approximation problem
[…]
$\beta_{*}^{(k)}=\operatorname*{argmin}_{\beta\in\mathbb{R}}\zeta_{\beta}^{2}\gamma_{\beta}^{2}.$ (4.1)
Recalling $r_{j}=|f_{j}-\xi^{(k)}(x_{j})|$ and
${\cal I}_{\widetilde{\boldsymbol{w}}}=\{j|\widetilde{w}_{j}>0,~{1\leq j\leq m}\}=\{j|w_{j}r_{j}>0,~{1\leq j\leq m}\}$
in Proposition 2.3, we have
$\nu(\beta):=\zeta_{\beta}^{2}\gamma_{\beta}^{2}=\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}\right).$ (4.2)
Proposition 4.1.
The function $\nu(\beta)$ given in (4.2) is convex.
Proof.
First, note that the derivative of $\nu(\beta)$ is
$\nu^{\prime}(\beta)=(\zeta_{\beta}^{2}\gamma_{\beta}^{2})^{\prime}$ $=-\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}\log r_{j}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}\right)$
$+\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}\log r_{j}\right).$
Moreover,
$\nu^{\prime\prime}(\beta)$
$=$ $\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}(\log r_{j})^{2}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}\right)+\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}(\log r_{j})^{2}\right)$
$-2\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}\log r_{j}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}\log r_{j}\right)$
$\geq$ $2\sqrt{\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}(\log r_{j})^{2}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}(\log r_{j})^{2}\right)}$
$-2\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}|\log r_{j}|\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}|\log r_{j}|\right)$
$\geq$ $2{\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}\left|q_{j}\right|^{2}r_{j}^{2-\beta}|\log r_{j}|\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}|\log r_{j}|\right)}$ (4.3)
$-2\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}|\log r_{j}|\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}|\log r_{j}|\right)$
$=$ $0,$
where the first inequality follows by using $a^{2}+b^{2}\geq 2{|ab|}$, while the second is due to $\|\boldsymbol{u}\|_{2}\|\boldsymbol{v}\|_{2}\geq|\boldsymbol{u}^{\operatorname{T}}\boldsymbol{v}|$. This implies that $\nu(\beta)$ is convex. ∎
Proposition 4.2.
For the polynomial minimax approximation problem, we have
  • (i)
    $\beta=1$ is the global minimizer of (4.1). In this sense, $\beta=1$ achieves the maximum of the lower bound in (3.2) and can be viewed as the near-optimal Lawson exponent in Lawson’s iteration;
  • (ii)
    for any $\beta\in[0,2]$, the sequence of Lawson’s iteration satisfies $d_{2}(\boldsymbol{w}^{(k+1)})\geq d_{2}(\boldsymbol{w}^{(k)})$.
Proof.
For this special case of $\beta=1$ and $q\equiv 1$, we only need to notice
$\nu^{\prime}(1)$ $=-\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}\log r_{j}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}\right)+\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}\log r_{j}\right)=0,$
and thus the conclusion (i) follows from Proposition 4.1.
For (ii), by the convexity of $\nu(\beta)$ (i.e., $\nu^{\prime\prime}(\beta)\geq 0$) and
${{d_{2}(\boldsymbol{w})}}=\nu(0)=\nu(2)\geq\nu(1)=\min_{\beta\in\mathbb{R}}\nu(\beta),$
[…]
Proposition 4.2 establishes the monotonic convergence and also locally the near-optimal choice of the Lawson exponent at each iteration. When $\beta=1$, the linear convergence of Lawson’s iteration was established in [6, 12]. Additionally, remedies for cases where assumption (A2) fails were discussed in [33, Chapter 13], while extensions of Lawson’s iteration for computing a minimax approximation in a general functional subspace spanned by a Chebyshev set were proposed in [34].
5 Monotonic convergence and complementary slackness for the rational minimax approximation problem
We now consider the convergence of d-Lawson for the rational minimax approximation problem. Different from the linear case, a difficulty arises from the numerator $(\boldsymbol{q}^{(k+1)})^{\operatorname{H}}W^{(k)}\boldsymbol{q}^{(k)}$ of the lower bound (3.2), in which $\boldsymbol{q}^{(k+1)}$ is also dependent on the Lawson exponent $\beta$. Explicitly expressing $\boldsymbol{q}^{(k+1)}$ in terms of $\beta$ in general is impossible because $\boldsymbol{q}^{(k+1)}$ is related with an eigenvector of the matrix pencil $(A_{\boldsymbol{w}},B_{\boldsymbol{w}})$. However, locally around $\beta=0$, it is possible to analyze the term $(\boldsymbol{q}^{(k+1)})^{\operatorname{H}}W^{(k)}\boldsymbol{q}^{(k)}$, and therefore, the lower bound in (3.2) with respect to $\beta$. Based on this observation, we can conclude that, generically, for any sufficiently small $\beta$, the monotonic property $d_{2}(\boldsymbol{w}^{(k+1)})\geq d_{2}(\boldsymbol{w}^{(k)})$ holds. This convergence result is consistent with the numerical experiments of the AAA-Lawson iteration where it is observed [14] that “taking $\beta$ to be smaller makes the algorithm much more robust”.
To develop our convergence, we need the following lemma.
Lemma 5.1.
Given $\boldsymbol{s}=[s_{1},\dots,s_{m}]^{\operatorname{T}}\in{\cal S}$, let $t(x)$ and $h(x)$ be strictly monotonically increasing on the interval $(a,b)$. Then for any $m$ points $x_{j}\in(a,b),~1\leq j\leq m$, we have
$\left(\sum_{j=1}^{m}s_{j}t(x_{j})\right)\left(\sum_{j=1}^{m}s_{j}h(x_{j})\right)\leq\sum_{j=1}^{m}s_{j}t(x_{j})h(x_{j});$ (5.1)
moreover, the equality in (5.1) holds if and only if $s_{i}s_{j}(x_{i}-x_{j})=0$ for all $1\leq i,j\leq m$.
Proof.
The result relies on the following inequality:
$t(x_{i})h(x_{j})+t(x_{j})h(x_{i})\leq t(x_{i})h(x_{i})+t(x_{j})h(x_{j}),~~\forall 1\leq i,j\leq m.$
In fact, the above is equivalent to $(t(x_{i})-t(x_{j}))(h(x_{j})-h(x_{i}))\leq 0$ which is true by assumptions on $t(x)$ and $h(x)$. The equality holds if and only if $x_{j}=x_{i}$. Thus, we have
$\left(\sum_{j=1}^{m}s_{j}t(x_{j})\right)\left(\sum_{j=1}^{m}s_{j}h(x_{j})\right)$
$=$ $\sum_{j=1}^{m}s_{j}^{2}t(x_{j})h(x_{j})+\sum_{1\leq i<j\leq n}s_{i}s_{j}\left(t(x_{i})h(x_{j})+t(x_{j})h(x_{i})\right)$
$\leq$ $\sum_{j=1}^{m}s_{j}^{2}t(x_{j})h(x_{j})+\sum_{1\leq i<j\leq n}s_{i}s_{j}\left(t(x_{i})h(x_{i})+t(x_{j})h(x_{j})\right)$
$=$ $\sum_{j=1}^{m}s_{j}(s_{1}+\dots+s_{m})t(x_{j})h(x_{j})=\sum_{j=1}^{m}s_{j}t(x_{j})h(x_{j}),$
and the equality holds if and only if $s_{i}s_{j}(x_{i}-x_{j})=0$ for all $1\leq i,j\leq m$. ∎
Theorem 5.1.
At the $k$th step of d-Lawson (Algorithm 1), for $\boldsymbol{w}^{(k)}\in{\cal S}$, assume $d_{2}(\boldsymbol{w}^{(k)})$ is a simple eigenvalue of the matrix pencil $(A_{\boldsymbol{w}^{(k)}},B_{\boldsymbol{w}^{(k)}})$ given in (2.14), and (A1) and (A2) hold. Suppose $r_{j}^{(k)}:=|f_{j}-\xi^{(k)}(x_{j})|>0,~\forall j\in{\cal I}_{\boldsymbol{w}^{(k)}}:=\{j|w_{j}^{(k)}\neq 0,~{1\leq j\leq m}\}$. Then
  • (i)
    there is a $\beta_{0}>0$ so that for any $\beta\in(0,\beta_{0})$, d-Lawson gives $d_{2}(\boldsymbol{w}^{(k+1)})\geq d_{2}(\boldsymbol{w}^{(k)})$;
  • (ii)
    for any sufficiently small $\beta>0$,
    $d_{2}(\boldsymbol{w}^{(k+1)})=d_{2}(\boldsymbol{w}^{(k)})\Longrightarrow w_{j}^{(k)}r_{j}^{(k)}\left(r_{j}^{(k)}-c\right)=0,~\forall 1\leq j\leq m,$ (5.2)
    where ${c=\sqrt{d_{2}(\boldsymbol{w}^{(k)})}}\leq\|\boldsymbol{f}-\xi^{(k)}(\boldsymbol{x})\|_{\infty};$
  • (iii)
    for the item (ii), if additionally, $c\geq\max_{j\not\in{\cal I}_{\boldsymbol{w}^{(k)}}}r_{j}^{(k)}$ and $\xi^{(k)}$ is irreducible, then $\xi^{(k)}$ is the minimax approximant of (1.1).
Proof.
For simplicity, we adopt the notation in the proof of Theorem 3.1 by omitting the superscripts, and we assume that $d_{2}(\boldsymbol{w})>0$ (the conclusions are trivial when $d_{2}(\boldsymbol{w})=0$ as $w_{j}r_{j}=0,~\forall j=1,2,\dots,m$). The idea for the proof is to express and estimate the lower bound in (3.2) using the real parameter $\beta$ around $\beta=0$.
Define $W(\beta)={\rm diag}(w_{1}(\beta),\dots,w_{m}(\beta))$ with
$w_{j}(\beta)=\frac{w_{j}r_{j}^{\beta}}{\gamma_{\beta}}=\frac{w_{j}|f_{j}-\xi(x_{j})|^{\beta}}{\sum_{j=1}^{m}w_{j}|f_{j}-\xi(x_{j})|^{\beta}}$
for which we have $w_{j}(0)=w_{j}$ and
$w_{j}^{\prime}(0)=w_{j}\log r_{j}-w_{j}\sum_{i=1}^{m}w_{i}\log r_{i},\quad\forall j\in{\cal I}_{\widetilde{\boldsymbol{w}}}$.
In (iii) of Proposition 2.1, based on assumptions (A1) and (A2), it is true that the matrices $Q_{p}(\beta),R_{p}(\beta),Q_{q}(\beta)$ and $R_{q}(\beta)$ in QR factorizations of $\sqrt{W(\beta)}\Phi=Q_{q}(\beta)R_{q}(\beta)$ and $\sqrt{W(\beta)}\Psi=Q_{p}(\beta)R_{p}(\beta)$, are all locally differentiable for sufficiently small $\beta$. Moreover, as $d_{2}(\boldsymbol{w}(0))$ is a simple eigenvalue of the matrix pencil $(A_{\boldsymbol{w}(0)},B_{\boldsymbol{w}(0)})$ given in (2.14), by (2.15), it is also a simple eigenvalue of the Hermitian matrix $S_{F}(\beta)-S_{qp}(\beta)S_{qp}(\beta)^{\operatorname{H}}$ at $\beta=0$. Thus, the continuity of eigenvalues implies that the smallest eigenvalue $d_{2}(\boldsymbol{w}(\beta))$ is a simple eigenvalue of the Hermitian $S_{F}(\beta)-S_{qp}(\beta)S_{qp}(\beta)^{\operatorname{H}}$ for any sufficiently small $\beta\in\mathbb{R}$. Alternatively, we can say that the eigenspace spanned by the unit-norm eigenvector $R_{q}(\beta)\boldsymbol{b}(\boldsymbol{w}(\beta))$ of $S_{F}(\beta)-S_{qp}(\beta)S_{qp}(\beta)^{\operatorname{H}}$ corresponding to $d_{2}(\boldsymbol{w}(\beta))$ is one-dimensional. Based on [26, Chapter 2.6.2], there is a continuously differentiable normalized eigenvector $R_{q}(\beta)\boldsymbol{b}(\boldsymbol{w}(\beta))$ with respect to $\beta$ around $\beta=0$. Moreover, noting that $\boldsymbol{q}(\beta)=\Phi\boldsymbol{b}(\beta)=\Phi(R_{q}(\beta))^{-1}R_{q}(\beta)\boldsymbol{b}(\beta)$, and also that $|\boldsymbol{q}(\beta)^{\operatorname{H}}W(0)\boldsymbol{q}(0)|$ does not change for different choices of a unit-norm eigenvector $R_{q}(\beta)\boldsymbol{b}(\boldsymbol{w}(\beta))$, in the following discussion, we can assume $\boldsymbol{q}(\beta)$ is continuously differentiable with respect to $\beta$ around $\beta=0$. Hence,
$\boldsymbol{q}(\beta)=\boldsymbol{q}(0)+\beta\boldsymbol{q}^{\prime}(0)+O(\beta^{2}).$
Since
$1=\boldsymbol{q}(\beta)^{\operatorname{H}}W(\beta)\boldsymbol{q}(\beta)\Longrightarrow{\rm Re}(\boldsymbol{q}^{\prime}(0)^{\operatorname{H}}W(0)\boldsymbol{q}(0))=-\frac{1}{2}\boldsymbol{q}(0)^{\operatorname{H}}W^{\prime}(0)\boldsymbol{q}(0),$
for any sufficiently small $\beta\in\mathbb{R}$, we have
$\left|\boldsymbol{q}(\beta)^{\operatorname{H}}W(0)\boldsymbol{q}(0)\right|$
$\geq\left|{\rm Re}(\boldsymbol{q}(\beta)^{\operatorname{H}}W(0)\boldsymbol{q}(0))\right|$
$=\left|\boldsymbol{q}(0)^{\operatorname{H}}W(0)\boldsymbol{q}(0)+\beta{\rm Re}(\boldsymbol{q}^{\prime}(0)^{\operatorname{H}}W(0)\boldsymbol{q}(0))\right|+O(\beta^{2})$
$=\left|1+\beta{\rm Re}(\boldsymbol{q}^{\prime}(0)^{\operatorname{H}}W(0)\boldsymbol{q}(0))\right|+O(\beta^{2})$
$=\left|1-\frac{\beta}{2}\boldsymbol{q}(0)^{\operatorname{H}}W^{\prime}(0)\boldsymbol{q}(0)\right|+O(\beta^{2})$
$=1-\frac{\beta}{2}\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}|q_{j}|^{2}w_{j}\log r_{j}-\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}|q_{j}|^{2}w_{j}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}\log r_{j}\right)\right)+O(\beta^{2})$
$=1-\frac{\beta}{2}\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}|q_{j}|^{2}w_{j}\log r_{j}-\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}\log r_{j}\right)+O(\beta^{2}).$
With this, we can write the lower bound in (3.2) as
$\widehat{\ell}(\beta):$ $=\frac{|\boldsymbol{q}(\beta)^{\operatorname{H}}W(0)\boldsymbol{q}(0)|}{\gamma_{\beta}~\zeta_{\beta}}$
$\geq\underbrace{\frac{1-\frac{\beta}{2}\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}|q_{j}|^{2}w_{j}\log r_{j}-\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}\log r_{j}\right)}{\sqrt{\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2-\beta}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}r_{j}^{\beta}\right)}}}_{=:\ell(\beta)}+O(\beta^{2})$
$=:\ell(\beta)+O(\beta^{2})$
locally at $\beta=0$.
For $\ell(\beta)$, by calculation, we have (with $q_{j}=q_{j}(0),~r_{j}=r_{j}(0)$ and $w_{j}=w_{j}(0)$)
$\ell^{\prime}(0)$
$=$ $\frac{1}{2\sqrt{(d_{2}(\boldsymbol{w}(0)))^{3}}}\left[\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2}\log r_{j}-\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}r_{j}^{2}\right)\left(\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}\log r_{j}\right)\right]$
$\geq$ $0$
where the last inequality is obtained by applying Lemma 5.1 with $s_{j}=w_{j}|q_{j}|^{2}$, $x_{j}=r_{j}$, $t(x)=x^{2}$ and $h(x)=\log x$ on the interval $(0,\infty)$. Furthermore, if there is a pair $(i,j)$ so that $i\in{\cal I}_{\widetilde{\boldsymbol{w}}}$}, $j\in{\cal I}_{\widetilde{\boldsymbol{w}}}$} and $w_{i}w_{j}(r_{i}-r_{j})\neq 0$, then we have $\ell^{\prime}(0)>0$ by Lemma 5.1. In that case, we know that there is a $\beta_{0}>0$ such that $\ell^{\prime}(\beta)>\frac{1}{2}\ell^{\prime}(0)$ and $\frac{\beta}{2}\ell^{\prime}(0)+O(\beta^{2})>0$ for any $\beta\in(0,\beta_{0})$, implying
$\widehat{\ell}(\beta)=\widehat{\ell}(0)+\int_{0}^{\beta}\widehat{\ell}^{\prime}(t)dt+O(\beta^{2})\geq\widehat{\ell}(0)+\frac{\beta}{2}\ell^{\prime}(0)+O(\beta^{2})>\widehat{\ell}(0).$
This shows that, if there is a pair $(i,j)$ so that $i,j\in{\cal I}_{\widetilde{\boldsymbol{w}}}$} and $w_{i}w_{j}(r_{i}-r_{j})\neq 0$, then a sufficiently small $\beta>0$ leads to $d_{2}(\boldsymbol{w}(\beta))>d_{2}(\boldsymbol{w}(0))$; conversely, for a sufficiently small $\beta>0$,
$d_{2}(\boldsymbol{w}(\beta))=d_{2}(\boldsymbol{w}(0))$ $\Longrightarrow w_{j}w_{i}(r_{i}-r_{j})=0,~\forall{i,j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}$
$\Longrightarrow r_{j}={\rm a~constant~}c,~\forall{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}$. (5.3)
Observe that $c\leq\max_{1\leq j\leq m}r_{j}=\|\boldsymbol{f}-\xi(\boldsymbol{x})\|_{\infty}$ and also
$d_{2}(\boldsymbol{w}(0))=\sum_{j=1}^{m}w_{j}|q_{j}r_{j}|^{2}=\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|r_{j}|^{2}|q_{j}|^{2}=c^{2}\sum_{j\in{\cal I}_{\widetilde{\boldsymbol{w}}}}w_{j}|q_{j}|^{2}{=}c^{2}\sum_{j=1}^{m}w_{j}|q_{j}|^{2}=c^{2},$ (5.4)
giving ${\sqrt{d_{2}(\boldsymbol{w}(0))}=c}\leq\|\boldsymbol{f}-\xi(\boldsymbol{x})\|_{\infty}.$ This proves items (i) and (ii).
For (iii), if we additionally have $c\geq\max_{j\not\in{\cal I}_{\boldsymbol{w}}}r_{j}$, then $c=\sqrt{d_{2}(\boldsymbol{w}(0))}=\|\boldsymbol{f}-\xi(\boldsymbol{x})\|_{\infty}$, which, according to Theorem 2.4 ([54, Theorem 4.3]), implies that Ruttan’s sufficient condition (or, equivalently, strong duality) for the minimax approximation problem is satisfied, and $\xi^{(k)}$ is the minimax approximant of (1.2). In this case, the result in (5.2) can be written as
$w_{j}\left(r_{j}-\|\boldsymbol{f}-\xi(\boldsymbol{x})\|_{\infty}\right)=0,~{1\leq j\leq m},$
which is the complementary slackness property in (2.13). ∎
Remark 5.1.
We remark first that the conclusion $w_{j}^{(k)}r_{j}^{(k)}\left(r_{j}^{(k)}-c\right)=0~{\forall j=1,2,\dots,m}$ in (5.2) is a certain complementary slackness. Indeed, it says that for node $x_{j}$ with $w_{j}^{(k)}>0$, either $r_{j}^{(k)}=0$ or $r_{j}^{(k)}=c$; that is, any node associated with positive weight either is an interpolation point or has the error $c$.
According to our proof for Theorem 5.1, at every iteration, the Lawson exponent $\beta_{0}>0$ in Theorem 5.1 is dependent on how far the continuously differentiable normalized eigenvector $R_{q}(\beta)\boldsymbol{b}(\boldsymbol{w}(\beta))$ of the Hermitian matrix $S_{F}(\beta)-S_{qp}(\beta)S_{qp}(\beta)^{\operatorname{H}}$ can be extended from $\beta=0$. Explicit formulation for $\beta_{0}$ is hard, but intuitively, the larger the gap between the smallest eigenvalue $d_{2}(\boldsymbol{w})$ and the next eigenvalue is, the larger $\beta_{0}$ is. Currently, the proof does not guarantee $\beta_{0}=1$, which, nevertheless, performs always well and is a recommended value in practice [54].
6 Conclusions
In this paper, we have established theoretical guarantees for Lawson’s iteration in solving both the linear and rational minimax approximation problems. For the rational minimax approximation problem, our results indicate that, generically, a small Lawson exponent $\beta>0$ leads to the monotonic convergence, and also reveal some interesting properties of d-Lawson. These theoretical guarantees, on the one hand, explain some numerical behaviors (for example, “taking $\beta$ to be smaller makes the algorithm much more robust[14]), and on the other hand, provide more insights on why and how Lawson’s updating scheme works. These theoretical results lay a solid ground for this version d-Lawson of Lawson’s iteration.
Acknowledgments
We express our gratitude to the anonymous referees for their comments and questions, which have significantly contributed to enhancing our manuscript in several ways: 1) motivating the addition of a new subsection 2.5 to more effectively elucidate intricate technical details and clarify the novelty of our theoretical analysis; 2) prompting us to conduct a more in-depth reflection on the relationship between (1.1) and (2.5); and 3) improving the overall presentation and clarity of the paper. We also thank Ren-Cang Li for the discussions and helpful suggestions, including a general version of Lemma 5.1.